### v1

# Convex Approximation of Probabilistic Reachable Sets from Small Samples Using Self-supervised Neural Networks

Abstract
Probabilistic Reachable Set (PRS) plays a crucial role in many fields of autonomous systems, yet efficiently generating PRS remains a significant challenge. This paper presents a learning approach to generating 2-dimensional PRS for states in a dynamic system. Traditional methods such as Hamilton-Jacobi reachability analysis, Monte Carlo, and Gaussian process classification face significant computational challenges or require detailed dynamics information, limiting their applicability in realistic situations. Existing data-driven methods may lack accuracy. To overcome these limitations, we propose leveraging neural networks, commonly used in imitation learning and computer vision, to imitate expert methods to generate PRS approximations. We trained the neural networks using a multi-label, self-supervised learning approach. We selected the fine-tuned convex approximation method as the expert to create expert PRS. Additionally, we continued sampling from the distribution to obtain a diverse array of sample sets. Given a small sample set, the trained neural networks can replicate the PRS approximation generated by the expert method, while the generation speed is much faster.
Note to PractitionersReachability analysis is one of the popular methods for solving real-time motion planning problems for dynamic systems under uncertainty while ensuring safety. However, determining the reachable set is challenging. Traditional methods are computationally expensive and often require extensive system information. Meanwhile, neural networks have been shown to construct distributions effectively. Hence, we propose leveraging the capability of neural networks to model the distribution of reachable sets. To simplify the creation of training datasets, we incorporate self-supervised learning in training the neural network. The proposed method offers several advantages: it is easy to use, generates accurate results quickly, and requires less system information. However, as neural networks are statistical methods relying on data-driven insights, their effectiveness inherently depends on the availability of sufficient historical data for the system.
Index Terms:
Probabilistic reachable set, Convex approximation, Deep Learning
I Introduction
Autonomous system operations face various uncertainties and hazards [1]. The reachable set approach is an important method in collision avoidance [2, 3], safety verification [4] and path planning [5] for Autonomous systems. However, finding the deterministic reachable sets with 100% confidence level is difficult [6]. Meanwhile, the deterministic reachable sets may be too large and over-conservative to be applicable [7]. Finding the Probabilistic Reachable Set (PRS) is more realistic, incorporating uncertainties and stochastic elements in the system dynamics or initial conditions [8]. A brief visualization of PRS is shown in Fig. 1, where PRS with different confidence levels are colored.
Figure 1: Probabilistic Reachable set represents the set of possible states a dynamical system can reach from an initial state. For instance, a PRS with a 95% confidence level means that 95% of the possible system trajectories, considering uncertainties or noise, will remain within the set bounds
I-A Review
Finding PRS remains challenging. One method for determining reachable sets under bounded uncertainties is the Hamilton-Jacobi (HJ) reachability analysis. This approach formulates the problem as a partial differential equation that can be addressed using numerical techniques. Although this method is manageable in scenarios with fewer dimensions, its efficiency significantly diminishes as the dimensionality increases [9][10]. Additionally, HJ reachability analysis requires detailed information about the dynamics or uncertainties. Meanwhile, strong assumptions are required, such as the absence of time correlations in parameter uncertainties. These limitations make it unsuitable for certain realistic situations [9].
Fortunately, we can identify the reachable sets using a learning-based method given a large dataset from an unknown distribution. Generating a reachable set with machine learning/data-driven has been explored. A trained locally-weighted regression algorithm and nonlinear SVM classifiers can predict the reachability of the two-point boundary value problem (2PBVP) [11]. Other data-driven methods such as the Gauss process [12], Monte Carlo [13], and stochastic disturbance [14] have also been developed to approximate reachable set. However, all of those methods have significant limitations in practice for three aspects: a) they all require a lot of samples to perform; b) they cannot provide fast and accurate probabilistic reachable set with adjustable confidence levels; c) their PRS bound is not convex and linear.
Deep learning technology has seen significant advancements recently in computer vision [15] and nature language process [16]. And those methods are further adopted in many different fields such as finance [17], cyber security [18], healthcare [19, 20] and robotics [21, 22, 23, 24]. Deep learning-based methods have been effectively applied to solve obstacle avoidance [25] and path planning problems [26, 27], demonstrating their suitability for addressing challenges in autonomous control tasks. The Universal Approximation Theorem [28], guarantees that a feedforward neural network can approximate any nonlinear real-valued function. The advantages of neural networks include their ability to generate complex outputs from simple inputs, as seen in generative models. Previous research shows that neural networks can work when massive data is massively missing [29] and can determine uncertainty [30]. Neural networks can rapidly extract the texture information of pavement from just a few images [31]. Neural networks indeed offer significant advantages in terms of reduced computing time. Also, the parallel nature of neural networks can lead to significant reductions in computing time [32]. Furthermore, Transformers [16] and self-supervised learning have revolutionized the field of artificial intelligence [33] and have become essential tools for developing more efficient, and versatile AI models when limited sample is available [34]. Therefore, we propose designing a transformer-based neural network to generate PRS, and we further employ self-supervised learning techniques to train this network.
I-B Contribution
In this paper, we propose a method that uses self-supervised and multi-label training to train a transformer-based neural network for generating convex approximations of PRS from a small number of samples. The numerical experiments prove that the neural network can generate reliable convex approximations when the given samples come from a distribution that has been learned by the network. This effective method relies on just one assumption:
  • Assumption 1: a large number of previous samples have been observed before the predicting task.
This assumption is satisfied in many fields such as autonomous driving [35], flight landing prediction [36], and robotics [37]. Our method offers three unique advantages that set it apart from other methods:
  • Advantage 1: No information about the desired system is needed; only a few samples of the desired distribution are required for PRS generating.
  • Advantage 2: The generation speed is extremely fast, making it suitable for online applications.
  • Advantage 3: All the PRS bound generated are convex and linear.
  • Advantage 4: Very few labeled data are needed for training, which simplifies the training process and makes large-scale training feasible.
II Problem Modeling
The main goal of the paper is to generate a convex approximation that can capture the PRS of an unknown static uncertainty whose data samples are gradually observed.
In this section, we will introduce and discuss the concepts of reachable sets and PRS associated with a dynamic system with uncertain discrete-time characteristics  [38, 39].
Let us consider a general discrete-time dynamic system expressed as
$\boldsymbol{s}_{k+1}=\boldsymbol{f}\left(\boldsymbol{s}_{k},\boldsymbol{u}_{k},\boldsymbol{\theta}_{k},\boldsymbol{w}_{k}\right),$ (1)
where, $\boldsymbol{s}_{k}\in\mathbb{R}^{2}$ denotes the state of the system (we only consider 2d dynamic systems in this paper, we use $s$ instead of $x$ like many other dynamic system research to distinguish x-dimension) at time step $k\in\{1,\ldots,T\}$, $\boldsymbol{u}_{k}\in\mathcal{U}_{k}$ denotes an uncertain control input, $\boldsymbol{\theta}_{k}\in\Theta$ denotes an uncertain parameter, $\boldsymbol{w}_{k}\in\mathbb{W}$ denotes an uncertain disturbance, and $s_{0}\in\mathcal{S}_{0}$ denotes the initial state. Diverging from most previous studies [9], this paper allows the sets $\mathcal{U}_{k}\subset\mathbb{R}^{m}$, $\Theta\subset\mathbb{R}^{p}$, $\mathbb{W}\subset\mathbb{R}^{q}$, and $\mathcal{S}_{0}\subset\mathbb{R}^{2}$ to be unbounded. Taking into account the uncertainties in $\boldsymbol{s}_{k}$, $\boldsymbol{u}_{k}$, $\boldsymbol{\theta}_{k}$, and $\boldsymbol{w}_{k}$, it follows that $\boldsymbol{x}_{k+1}$ becomes a random vector following an unknown probability distribution. Then at time $T$, the reachable set $\mathcal{S}_{T}$ of the dynamic system in Eq. 1 is defined to be
[…]
$\boldsymbol{s}_{0}\in\mathcal{S}_{0},\boldsymbol{u}_{k-1}\in\mathcal{U}_{k-1},\boldsymbol{\theta}_{k-1}\in\Theta,\boldsymbol{w}_{k-1}\in\mathbb{W},$
$k\in\{1,\ldots,T\}\},$
where $\boldsymbol{s}_{0}$ denotes the initial state and $\mathcal{S}_{0}$ denotes the initial set. Observe that because the sets $\mathcal{U}_{k-1}$, $\Theta$, $\mathbb{W}$, and $\mathcal{S}_{0}$ can be unbounded, the resulting reachable set $\mathcal{S}_{T}$ could also be unbounded. This implies that determining a bounded set where the state is definitively contained might not be feasible. Therefore, our goal is to generate a bounded set where the probability of the state being within this set exceeds a certain confidence level, which is called a Probabilistic Reachable Set (PRS). At time $k$, a bounded set $\tilde{\mathcal{S}}_{k}$ can be qualified as a PRS of the dynamic system defined in Eq. 1 at confidence level $\alpha$ if and only if
$\operatorname{Pr}\left(\boldsymbol{s}_{k}\in\tilde{\mathcal{S}}_{k}\right)\geq\alpha.$ (3)
It indicates that the probability of the system reaching the PRS at time $k$ exceeds $\alpha$. For example, if $\alpha=95\%$ and the system runs an infinite number of times, the system will be outside the PRS less than 5% of the time. When $\alpha=100\%$ and a PRS $\tilde{\mathcal{S}}_{k}$ does exist, then the reachable set $\mathcal{S}_{k}$ defined in Eq. 2 belongs to the $\tilde{\mathcal{S}}_{k}$.
The previous expert method [40] can find multiple $\tilde{\mathcal{S}}_{k}$ and all the founded $\tilde{\mathcal{S}}_{k}$ are convex and as tight as possible. We define the set $\tilde{\mathcal{S}}_{k}$ using the positions of the four vertices, represented as $\tilde{\mathcal{V}}_{k}\in\mathbb{R}^{4\times 2}$. The trained neural network should be able to imitate the expert and generate a similar PRS that has a confidence level close to $\alpha$, as shown in Eq. 4,
$\tilde{\mathcal{V}}_{k}=NN(w,\mathbf{s_{k}})$ (4)
where $w$ is the trained weights of the neural network, $\mathbf{s_{k}}=\{s_{k}^{i}\}_{i=1}^{n}$ is a collection of samples from the real distribution of the reachable set with $s_{k}^{i}\sim\boldsymbol{s}_{k}$, and $n$ as the number of the samples. The desired $n$ should be smaller than 100.
In the rest of the paper, we only consider the reachable set for a specific single time step, for example, $k=T$. So to simplify the notation, we use $\mathcal{S}$ to refer the actual reachable set, $\tilde{\mathcal{S}}$ to refer the PRS, $\tilde{\mathcal{V}}$ as the four vertices of the PRS, and $\boldsymbol{s}$ as the state of the system, $\mathbf{s}$ as the collection of samples from the reachable set at desired time point.
III Dataset and Expert Method
(a) Example of 2d-binominal distribution and sample expert convex PRS
(b) Example of fan-shape distribution and sample expert convex PRS
(c) Example of GMR trajectory predictor distribution and sample expert convex PRS
Figure 2: Distribution and convex PRS demo
III-A Target distribution
In this paper, we evaluate the proposed method to determine if it can approximate three different probabilistic distributions, all of which are based on the Gaussian distribution, commonly used in aircraft trajectory modeling [41]. Any sample from the target probabilistic distributions is 2-dimensional data, $\mathbf{s}\in R^{2}$.
III-A1 2d-binomial distribution
2d-binomial distribution [38] forms clusters around two different coordinate points. The sampling plot of the 2d-binominal distribution is shown in Fig.2(a). The sample process includes two steps, first is randomly selecting one of two center points, second is adding Gaussian noise to both the x- and y-coordinates of the chosen center point. Let $\boldsymbol{C}_{1}=(x_{1},y_{1})$ and $\boldsymbol{C}_{2}=(x_{2},y_{2})$ represent the two center points. Let $\mathbf{N}=\begin{pmatrix}N_{x}\\ N_{y}\end{pmatrix}$ be the Gaussian noise added to both coordinates, where $N_{x}\sim\mathcal{N}(0,\sigma^{2})$ and $N_{y}\sim\mathcal{N}(0,\sigma^{2})$ are independent Gaussian random variables with mean 0 and variance $\sigma^{2}$. The sample of the 2d-binomial distribution will be $\boldsymbol{s}=\boldsymbol{C}_{i}+\mathbf{N},i\in\{1,2\}$. The probability density function(PDF) of the 2d-binomial distribution is easy to calculate:
$P(\boldsymbol{s})=\frac{1}{2}\cdot\mathcal{N}(\boldsymbol{s}|\boldsymbol{C}_{1},\sigma^{2}\mathbf{I})+\frac{1}{2}\cdot\mathcal{N}(\boldsymbol{s}|\boldsymbol{C}_{2},\sigma^{2}\mathbf{I}),$ (5)
where, $\mathcal{N}(\boldsymbol{s}|\boldsymbol{C}_{i},\sigma^{2}\mathbf{I})$ represents the bivariate Gaussian distribution centered at $\boldsymbol{C}_{i}=(x_{i},y_{i})$ with covariance matrix $\sigma^{2}\mathbf{I}$ (identity matrix scaled by $\sigma^{2}$, meaning independent noise with the same variance for $x$ and $y$).
III-A2 Fanshape distribution
The fan-shape distribution[38] represents points sampled from a distribution of motion paths influenced by truncated normal distributions of speed and direction, capturing realistic variations in both parameters. The sampling plot of the fanshape distribution is shown in Fig.2(b).
[…]
The position $\mathbf{s}:=[x_{k},y_{k}]^{\top}$ of the flight at time $k$ is modeled by the following discrete-time dynamic system:
$\begin{bmatrix}x_{k}\\ y_{k}\end{bmatrix}=\begin{bmatrix}x_{k-1}\\ y_{k-1}\end{bmatrix}+\begin{bmatrix}\cos\theta\\ \sin\theta\end{bmatrix}v\Delta t,$
where $[x_{k-1},y_{k-1}]^{\top}$ is the position of the fight at the previous time step $k-1$. $[x_{k},y_{k}]^{\top}$ is the position of the flight at the time step $k$. $v$ is the speed and $\theta$ is heading angle. Due to the uncertainties in $v$ and $\theta$, the position $\mathbf{s}$ is a random vector that obeys a non-Gaussian distribution, which is referred to as the fanshaped distribution. The PDF of the position is difficult to derive.
III-A3 GMR trajectory predictor
GMR trajectory predictor can regress the possibility distribution of the future position of flight near terminal [42] with multiple Gaussian distributions given the past trajectory. The sampling plot of the GMR trajectory predictor distribution is shown in Fig.2(c). The regressed conditional possibility distribution of the future position of flight is defined by the following equation:
$P(\boldsymbol{s}\mid\boldsymbol{s}^{past})=\sum_{k=1}^{K}\pi_{\boldsymbol{s}\mid\boldsymbol{s}^{past}_{k}}\mathcal{N}_{k}\left(\boldsymbol{s}\mid\boldsymbol{\mu}_{\boldsymbol{s}\mid\boldsymbol{s}^{past}_{k}},\boldsymbol{\Sigma}_{\boldsymbol{s}\mid\boldsymbol{s}^{past}_{k}}\right),$ (6)
where where $\mathcal{N}_{k}\left(\boldsymbol{s}\mid\boldsymbol{\mu}_{\boldsymbol{s}\mid\boldsymbol{s}^{past}_{k}},\boldsymbol{\Sigma}_{\boldsymbol{s}\mid\boldsymbol{s}^{past}_{k}}\right)$ are Gaussian distributions with mean $\boldsymbol{\mu}_{\boldsymbol{s}\mid\boldsymbol{s}^{past}_{k}}$ and covariance $\boldsymbol{\Sigma}_{\boldsymbol{s}\mid\boldsymbol{s}^{past}_{k}}$, $K$ is the number of Gaussians, and, $\boldsymbol{s}^{past}_{k}$ is the input and $\boldsymbol{s}$ is the output, where $\boldsymbol{s}^{past}_{k}$ is the combined past trajectory and future trajectory guide of the ego agent. $\boldsymbol{s}$ is the trajectory prediction. $\pi_{k}\in[0,1]$ are weights that sum up to one. If $\pi_{k}$ is larger, the more likely the $\boldsymbol{s}$ belongs to $\mathcal{N}_{k}$. We use three-time step prediction as the desired distribution.
III-B Expert method
The expert method is proposed in [40]. The expert method can generate the PRS of an uncertain dynamic system. This data-driven method builds on a kernel density estimator (KDE) that can generate the PRS, accelerated by a fast Fourier transform (FFT). These KDE values are then utilized within an Integer Linear Programming (ILP) framework to find a minimal parallelogram box that approximates the probabilistic geofence. This ILP solution effectively encapsulates the area where the state is expected to stay with high confidence.
A brief process of the ILP Heuristic Algorithm is shown in the Algorithm 1.
Algorithm 1 ILP Heuristic Algorithm
1:function BoxApprox($r_{\min},c_{\min},r_{\max},c_{\max},\omega_{ij}$)
2:     $\omega^{\prime}_{ij}\leftarrow\omega_{ij}.\text{copy}()$
3:     $z_{ij}[1:N][1:N]\leftarrow 0$
4:     $r_{\min},c_{\min}\leftarrow\arg\max_{i,j}\omega^{\prime}_{ij}$
5:     $r_{\max}\leftarrow r_{\min}$, $c_{\max}\leftarrow c_{\min}$
6:     while $\sum_{i=1}^{N}\sum_{j=1}^{N}\omega_{ij}z_{ij}<\alpha\sum_{i}\sum_{j}\omega_{ij}$ do
7:         $i,j\leftarrow\arg\max_{i,j}\omega^{\prime}_{ij}$
8:         $\omega^{\prime}_{ij}[i,j]\leftarrow 0.0$
9:         $r_{\min}\leftarrow\min(r_{\min},i)$, $r_{\max}\leftarrow\max(r_{\max},i)$
10:         $c_{\min}\leftarrow\min(c_{\min},j)$, $c_{\max}\leftarrow\max(c_{\max},j)$
11:         $z_{ij}[r_{\min}:r_{\max}][c_{\min}:c_{\max}]\leftarrow 1$
12:     end while
13:     return $z_{ij}$
14:end function
The ILP Heuristic Algorithm approximates solutions for integer linear programming problems using heuristics. It starts with a weight matrix $\omega_{ij}$ and a zero matrix $z_{ij}$, setting initial indices based on $\omega_{ij}$’s maximum values. The algorithm loops to update $z_{ij}$ until the product of $z_{ij}$ and $\omega_{ij}$ meets a threshold $\alpha$. Each iteration zeros out the maximum $\omega_{ij}$ and adjusts boundary indices. The corresponding $z_{ij}$ entry is set to one within these boundaries, yielding the matrix as an efficient approximation of the ILP solution. The expert convex generated is shown in Fig. 2. The black points represent the distribution and purple convex covers a desired percent of the distribution.
III-C Training dataset generation
Theoretically, there are unlimited correct convex approximations for PRS given a probability distribution. At the same time, the expert method introduced in the previous section can generate a very accurate convex approximation if enough samples are given. Therefore, for each distribution $P$ and each disired confidence level $\alpha$, we can generate a set of correct convex approximation $\mathbf{\tilde{\mathbf{S}}}_{p,\alpha}=\{\tilde{\mathcal{S}}_{P,\alpha}^{i}\}_{i=1}^{n}$. In this paper, the correct convex approximation has 1% error toleration. Therefore, we consider the approximation is correct if $\alpha+0.01\geq\operatorname{Pr}\left(\boldsymbol{s}\in\tilde{\mathcal{S}}_{\alpha}\right)\geq\alpha$. Each correct convex approximation $\tilde{\mathcal{S}}_{P,\alpha}$ is defined by four vertices $\tilde{\mathcal{V}}_{P,\alpha}$. Then the set of vertices of the correct convex approximation is denoted as $\mathbf{\tilde{\mathbf{V}}}_{p,\alpha}=\{\tilde{\mathcal{V}}_{P,\alpha}^{i}\}_{i=1}^{n}$. The usage of the generated correct convex approximation will be introduced in the method section below.
IV Method
IV-A Self-supervised learning
Figure 3: Masked distribution: we only show a few samples to the neural networks, and ask neural networks to generate the corresponding PRS.
Previous research [33] has proven that the transformer can process the visible (unmasked) parts of the data, reconstructing the entire input from the encoded, incomplete data (masked data). Therefore, if we only have a few samples from the distribution, the transformer may be able to reconstruct the entire distribution.
As the previous section III, we have the vertices of expert convex approximation $\mathbf{\tilde{\mathbf{V}}}_{p,\alpha}$. Fig.3 shows the examples of how we mask the distribution. During the training, we randomly mask most of the sample points, only leaving a few samples visible to the transformer, and ask the transformer to generate the corresponding PRS.
With this self-supervised learning strategy, we are able to show many sets of visible samples of one distribution to the neural networks. When we show the stranger samples to the trained neural networks, the trained neural network is able to identify the distribution by recognizing and relating the visible samples to the samples the neural networks have seen during the training process.
IV-B Multi-label learning and loss function
Figure 4: Example of multi-label learning
As introduced in section II, one distribution can have multiple correct convex approximations. Therefore, we can use multi-label learning instead of single-label learning. Multi-label learning not only reduces computational overhead but also allows the model to share learned features across multiple labels, potentially leading to better generalization. Fig. 4 shows an example of multi-label learning. The neural network generated an output based on the current weights of the neural network. Then the losses between each expert and the output are calculated and the smallest loss will be used to train the neural network. For the real training, we will have more expert PRS samples. Like many regression machine learning research studies, we use the MSE loss function as shown in Eq. 7. MSE loss can help the neural network minimize the average error between the expert and the output.
$\text{Loss}=\min_{j\in[1,10]}{\frac{1}{8}\sum_{i=1}^{8}(\hat{y}_{ij}-y_{i})^{2}}\quad\text{where}\quad\hat{y}_{ij}\in\mathbf{\tilde{\mathbf{V}}}_{p,\alpha}$ (7)
where:
  • we have 8 (four 2d vertices) elements in the output vector,
  • $y_{i}$ is is the generated value for the $i$-th element,
  • $\hat{y}_{ij}$ is the $j$-th expert vertices for the $i$-th element.
IV-C Neural networks architecture
The overall architecture of the neural network we used is shown in Fig. 5, which is designed based on our previous work [43]. The input to the neural network consists of sampled points and the desired confidence level. Before sending these inputs to the neural network, we normalize them first. The confidence level and the sample points are then embedded in a higher dimension using a linear layer. The embedded items are concatenated together. These concatenated embeddings are then sent to the transformer layer. In the transformer, the channel representing the confidence level interacts with all the point channels. Finally, we send the confidence level channel to an output layer, resulting in four vertices of the convex hull.
Figure 5: Neural network architecture
Figure 6: Concatenated tensor
IV-D Input normalization, embedding, and concatenate
Because sample points are drawn from different distributions, they have very different means and variances, which can cause challenges for the training process of neural networks. Normalizing those points before feeding them into the neural network is very crucial for several reasons. First, if certain inputs have larger values and higher variance than others, they may dominate the training process. For example, if one target (or feature) has a very large value compared to others, its gradient will be disproportionately large. Second, normalization can accelerate the training process. When the scales of the inputs are not consistent, the model might spend more time adjusting weights to accommodate the varying scales of input data rather than learning meaningful patterns. Third, normalization can prevent overfitting by reducing bias toward dominant features and indirect regularization effects.
We use the Min-Max scaling [44] to normalize the sample points. The Min-Max scaling equation is defined by:
$x^{\prime}=\frac{x-\min(x)}{\max(x)-\min(x)}.$ (8)
Min-max scaling scales sample points to the range [0, 1] without changing the relative position between points. At the same time, since all the points are sampled from a reasonable distribution, so no outlier can heavily influence the normalization.
After normalizing, we use a single linear layer to embed the inputs. Compared to vision or natural language inputs, the inputs in our task are much more straightforward to understand. We primarily aim to determine the relationship between points rather than to interpret their meanings. Additionally, the transformer model already incorporates non-linearities. Therefore, we choose to embed the inputs into relatively low dimensions and use only a single linear layer to decrease the computing complexity.
After embedding, the point tensor has dimensions ($N$, $C$), where $N$ is the dimension of the series of points and $C$ is commonly referred to as the channel in the field of machine learning. We concatenated the confidence level tenor and the point tensor together along the $N$ dimension. So, the concatenated tensor has dimensions ($N+1$, $C$), while the first channel is the channel of the confidence level. The concatenated tensor then looks like the Fig. 6.
IV-E Transformer
For this work, we only use the transformer encoder [16] since we output the vertex at once instead of one by one. The encoder is composed of a stack of 6 identical layers, each layer has two sub-layers, a multi-head self-attention mechanism, and the feedforward network. We also kept the residual connection and layer normalization. Compared to other popular layers, such as the recurrent and convolutional layers, self-attention has three main advantages. First, self-attention has lower computational complexity per layer. Second, it requires fewer sequential operations since most computations can be parallelized. The third advantage is that self-attention can learn the dependencies between two points that are far from each other. For example, if we use CNNs, we may input points patch by patch, then it is difficult to learn the relationship between the first patch and the last patch. In contrast, each point interacts with all the other points at the same time in self-attention.
IV-F Output
In the previous masked autoencoder research. Mask tokens are introduced after the encoder, so the decoder will know which missed position is required to be generated. However, unlike reconstructing assignments, we do not have mask tokens. We do not know the visible samples are from which part. In fact, unlike CV and NLP problems, the sequence does not matter in our problem. For example, ”The dog chased the cat” and ”The cat chased the dog” are two completely different sentences, even though they have exactly the same words. In contrast, in our problem, the point set [(1, 1), (2, 2)], and [(2,2), (1, 1)] have the same distribution. Therefore, no additional global context is required. Furthermore, the goal of the proposed neural network is to generate a convex bound instead of reconstructing the whole distribution. All in all, a simple output layer is enough to generate the convex bound with the features captured from the visible samples.
We use a multilayered output like ViT [15] instead of a single linear layer like the original transformer. The multi-layer output can further process the features extracted from the points channels and refine the final classification. We only used the transformed confidence level channel to output. The transformer allows the confidence level channel to attend to all other channels (embedded points), gathering contextual information from all the points. Finally, we need to denormalize the output, since the points we want to bound are normalized.
IV-G Parameters
For the transformer, embedding dimension $N_{emd}$ is 864. We use 6 attention heads, so each head processes 144 embedding dimensions of each channel. The number of attention layer is also 6. Optimization is performed using the Adam optimizer with a dynamic learning rate to ensure gradual convergence. Additionally, a dropout rate of 0.2 is applied to prevent overfitting by randomly omitting certain neurons during training.
V Result
(a) 2d-binominal distribution
(b) Fanshape distribution
(c) GMR trajectory predictor distribution
Figure 7: Example outputs if three methods take the same visible points, visible points are enlarged as yellow points for greater visibility.
V-A Experiment setting
We compare the proposed method with the ILP Heuristic (ILP), which is the expert method, and the bounding box (BB) method [45][46], a popular approach to finding a convex approximation of the PRS. Both baseline methods are KDE-based. Fig. 7 are examples of point distribution and convex hulls generated by different methods given the same visible points. All the experiences are run in Intel(R) Core(TM) i9-12900KF and NVIDIA GeForce RTX 3090.
V-A1 Training dataset
In each training iteration, we randomly assign ten sets of expert vertices as the target and randomly pick 100 samples (300 for the GMR trajectory predictor) from the desired distribution as the neural network’s input. To accelerate the training process, we prepare 100 sets of expert vertices for each desired confidence level and prepare 320,000 sets of samples before training begins. To make the required expert vertices, it will cost experts about 24 hours for all distributions. In addition to expert vertices generating, we also need to generate training input samples before training. The sample generation time varies across different distributions. For the binomial distribution, it takes only about 2 minutes to generate 320,000 sets of samples. For the fanshape distribution, the process takes approximately 40 minutes to produce the same number of sample sets. In the case of the GMR trajectory predictor, it requires around 30 hours to generate 320,000 sets of samples. Generally, we can prepare the training dataset for a known distribution within a few days. It supports that the Advantage 4 of the proposed method, the training dataset is easy to make. Additionally, the proposed method exhibits Advantage 3: all PRS are guaranteed to be convex and linear, as the neural network is specifically designed to generate only the vertices of convex sets.
V-A2 PRS generating
We tested the trained neural networks by presenting them with 5,000 random tasks (1000 for each $\alpha$) in each distribution. Each task contains 100 (300 for GMR) points of the distribution and has never been seen by the neural network. The neural network will generate the corresponding PRS for those points. At the same time, we also generate PRS with BB and ILP given the same points as the baseline. However, since BB and ILP are not learning-based and do not require a training process, comparing them directly may not be entirely fair. To adjust for this, we also compared the results using PRS generated by BB and ILP with 20 times more samples to thoroughly evaluate our proposed method. Besides the baseline method, we also run the MINLP Optimal Algorithm [47] to determine the smallest area of a convex satisfies the coverage requirement.
V-A3 PRS evaluating
To evaluate a PRS, we first need to determine the coverage. To represent the real distribution, we sampled 100,000 points from the desired distribution. The resulting coverage was determined by counting the percentage of these sampled points that fell within the generated PRS.
Mean Squared Error(MSE) is the metric we used to evaluate the accuracy with which the PRS approximates the desired confidence level:
$\text{MSE}=\frac{1}{n}\sum_{i=1}^{n}\left(\alpha_{i}-\hat{\alpha}_{i}\right)^{2}.$ (9)
$\alpha$ is the coverage of the generated PRS and $\hat{\alpha}_{i}$ is the desired coverage. n is 1000 because we test 1000 times for each $\alpha$. We also recorded the minimum and maximum coverage values to evaluate performance stability. While a low MSE indicates that the PRS is accurate, it does not guarantee correctness. Therefore, it is essential to compare the percentage of instances where the PRS exceeds the desired coverage. We called it the correctness rate.
V-A4 Generating time
Generating time is a crucial property in deciding whether a method can be used in online applications. Therefore, we compare the time to generate one correct PRS by each method.
V-B Compare coverage accuracy, correctness, and area to baseline
2d-binomial NN(our method) BB(100) ILP(100)
confidence level MSE(range) MSE(range) MSE(range)
80% 0.42%(79.12% - 82.05%) 11.04%(87.56% - 99.42%) 3.66%(72.00% - 93.54%)
85% 0.35%(84.45% - 86.63%) 10.18%(91.98% - 99.80%) 3.70%(72.17% - 95.76%)
90% 0.24%(89.51% - 91.01%) 8.20%(93.26% - 99.94%) 4.98%(86.16% - 98.78%)
95% 0.62%(94.08% - 96.29%) 4.20%(94.84% - 99.99%) 3.25%(92.45% - 99.73%)
99% 0.87%(97.66% - 99.94%) 1.87%(97.09% - 99.99%) 0.80%(98.97% - 99.99%)
TABLE I: MSE and range on 2d-binomial distribution distribution compare to baseline with a small number of sample
fanshape NN(our method) BB(100) ILP(100)
confidence level MSE(range) MSE(range) MSE(range)
80% 0.32%(78.95% - 80.44%) 11.98%(80.28% - 99.80%) 2.83%(75.04% - 91.87%)
85% 0.25%(84.44% - 86.21%) 9.88%(85.39% - 99.91%) 2.95%(75.24% - 94.01%)
90% 0.23%(89.40% - 90.91%) 7.36%(89.60% - 99.99%) 2.69%(75.63% - 98.43%)
95% 0.49%(95.22% - 96.02%) 4.32%(92.60% - 100%) 2.20%(85.67% - 99.88%)
99% 0.98%(99.96% - 99.99%) 1.33%(98.03% - 100%) 1.12%(97.86% - 100%)
TABLE II: MSE and range on fanshape distribution distribution compare to baseline with a small number of sample
GMR trajectory predictor NN(our method) BB(300) ILP(300)
confidence level MSE(range) MSE(range) MSE(range)
80% 1.79%(79.54% - 84.27%) 11.31%(84.04% - 96.41%) 2.74%(75.00% - 88.70%)
85% 1.76%(82.34% - 88.25%) 9.56%(88.09% - 98.02%) 2.62%(75.31% - 90.94%)
90% 1.13%(88.31% - 91.8%) 7.37%(93.11% - 99.40%) 1.76%(81.48% - 96.56%)
95% 0.85%(94.14\$ - 97.83%) 4.09%(95.82% - 99.95%) 1.11%(90.61% - 98.80%)
99% 0.58%(99.18% - 99.94%) 0.80%(98.42% - 100%) 0.40%(97.03% - 99.95%)
TABLE III: MSE and range on GMR trajectory predictor distribution compare to baseline with a small number of sample
2d-binomial NN(our method) BB(2000) ILP(2000)
confidence level MSE(range) MSE(range) MSE(range)
80% 0.42%(79.12% - 82.05%) 16.26%(93.24% - 98.43%) 4.12%(72.01% - 86.43%)
85% 0.35%(84.45% - 86.63%) 12.10%(95.65% - 98.49%) 3.15%(74.27% - 93.60%)
90% 0.24%(89.51% - 91.01%) 8.12%(96.52% - 99.63%) 4.37%(89.70% - 98.32%)
95% 0.62%(94.08% - 96.29%) 4.28%(98.64% - 99.86%) 3.22%(95.48% - 99.53%)
99% 0.87%(97.66% - 99.94%) 0.93%(99.73% - 99.99%) 0.86%(99.49% - 99.98%)
TABLE IV: MSE and range on 2d-binomial distribution distribution compare to baseline with a large number of sample
fanshape NN(our method) BB(2000) ILP(2000)
confidence level MSE(range) MSE(range) MSE(range)
80% 0.32%(78.95% - 80.44%) 12.73%(88.66% - 96.29%) 2.01%(75.36% - 84.46%)
85% 0.25%(84.44% - 86.21%) 10.98%(91.93% - 98.82%) 2.31%(81.43% - 90.18%)
90% 0.23%(89.40% - 90.91%) 8.63%(95.93% - 99.75%) 1.90%(87.40% - 94.16% )
95% 0.49%(95.22% - 96.02%) 4.81%(98.59% - 100%) 2.72%(94.35% - 99.06%)
99% 0.98%(99.96% - 99.99%) 0.99%(99.96% - 100%) 0.98%(99.19% - 100%)
TABLE V: MSE and range on fanshape distribution distribution compare to baseline with a large number of sample
GMR trajectory predictor NN(our method) BB(6000) ILP(6000)
confidence level MSE(range) MSE(range) MSE(range)
80% 1.79%(79.54% - 84.27%) 10.47%(85.60% - 93.33%) 2.40%(75.00% - 81.90%)
85% 1.76%(82.34% - 88.25%) 9.45%(90.69% - 96.35%) 3.01%(85.60% - 93.33%)
90% 1.13%(88.31% - 91.8%) 7.44%(95.89% - 98.67%) 1.37%(85.62% - 94.55%)
95% 0.85%(94.14\$ - 97.83%) 4.22%(97.98% - 99.67%) 0.74%(92.35% - 97.50%)
99% 0.58%(99.18% - 99.94%) 0.91%(99.72% - 99.97%) 0.25%(98.28% - 99.61%)
TABLE VI: MSE and range on GMR trajectory predictor distribution compare to baseline with a large number of sample
2d-binomial correctness rate
confidence level NN BB(100) BB(2000) ILP(100) ILP(2000) optimal(10000)
80% 99.7% 100% 100% 31.4% 14.7% 100%
85% 99.8% 100% 100% 67.3% 57.9% 100%
90% 99.8% 100% 100% 97.2% 99.8% 100%
95% 98.9% 99.9% 100% 99.1% 100% 100%
99% 98.7% 95.2 100% 99.9% 100% 100%
confidence level PRS area
80% 35.22 45.65 46.14 36.06 33.95 33.01
85% 41.58 52.16 50.12 42.42 40.69 34.23
90% 51.38 63.18 59.10 53.86 52.35 42.79
95% 64.53 75.85 71.90 67.46 66.08 55.01
99% 99.03 101.49 106.67 95.28 96.10 73.96
TABLE VII: correctness rate/PRS area for each desired confidence level on 2d-binomial distribution
fanshape correctness rate
confidence level NN BB(100) BB(2000) ILP(100) ILP(2000) optimal(10000)
80% 98.7% 100% 100% 57.7% 79.8% 100%
85% 99.8% 100% 100% 52.4% 64.4% 100%
90% 99.7% 99.8% 100% 52.8% 82.3% 100%
95% 100% 99.4% 100% 74.8% 98.9% 100%
99% 100% 61% 100% 59.7% 100% 100%
confidence level PRS area
80% 1415.07 2342.52 2309.52 1783.52 1514.04 1172.71
85% 1762.30 2790.63 2815.55 2138.71 1902.27 1579.52
90% 2229.01 3367.96 3470.37 2652.60 2365.27 2107.79
95% 3530.87 4525.46 4576.97 3382.33 3374.61 2742.49
99% 4833.31 5749.98 6127.58 4967.29 4820.67 3711.22
TABLE VIII: correctness rate/PRS area for each desired confidence level on fanshape distribution
GMR trajectory predictor correctness rate
confidence level NN BB(300) BB(6000) ILP(300) ILP(6000) optimal(30000)
80% 93.1% 100% 100% 10.8% 3.9% 100%
85% 95.2% 100% 100% 27.9% 16.2% 100%
90% 99.8% 100% 100% 69.2% 55.8% 100%
95% 95.6% 100% 100% 68.2% 65.4% 100%
99% 100% 99.7% 100% 56.5% 75.8% 100%
confidence level PRS area
80% 2.75E-3 3.64E-3 3.46E-3 2.59E-3 2.39E-3 2.15E-3
85% 3.44E-3 4.41E-3 4.35E-3 2.99E3 2.89E-3 2.39E-3
90% 4.41E-3 7.37E-3 5.81E-3 3.83E-3 3.67E-3 2.99E-3
95% 5.69E-3 7.89E-3 7.98E-3 5.03E-3 4.87E-3 3.95E-3
99% 9.21E-3 11.19E-3 11.16E-3 8.01E-3 7.63E-3 6.46E-3
TABLE IX: correctness rate/PRS area for each desired confidence level on GMR trajectory predictor distribution
First, we aim to evaluate whether the proposed method can effectively produce accurate PRS. Table III presents the MSE and range comparisons between the proposed method and two baseline approaches, given the exact same visible points on the 2d binomial distribution. Our method consistently achieves lower MSE and a narrower range than the baseline methods, except at a 99% confidence level, where its performance closely matches that of ILP. Table III shows the MSE and range results on the fanshape distribution. The proposed method demonstrates exceptional performance, with significantly lower MSE and a tighter range across all confidence levels. Finally, Table III corroborates the findings from the previous tables, further showcasing that the proposed method achieves substantially lower MSE and a consistently tighter range, underscoring its superior performance. All in all, we found out that our method can produce very accurate PRS. All MSE is lower than 1%, with smaller intervals, indicating our method is very stable and consistent. In contrast, KDE-based baseline methods have higher MSE and variance especially when the desired confidence level is lower. On the other hand, it is not entirely fair to compare KDE-based methods directly, as they do not require a training process. Theoretically, estimating a distribution accurately with a small number of samples is nearly impossible. Therefore, it is also necessary to evaluate whether the trained neural networks can outperform the baseline methods when a sufficient number of samples are provided. Tabel VI VI VI present the MSE and range comparisons between the proposed method (can have only 100 visible samples) and two baseline methods, given 2000 visible samples on three test distributions. While the results of the two baseline methods show slight improvements, they still fall significantly short of the performance achieved by our method.
Second, we want to ensure the convex bound generated by the neural networks is correct and useful. Table IX shows the performance of the neural network on 2d-binomial distribution. The neural network gets at least 98.7% correctness rate and only misses a few tests. Meanwhile, the average area we got is not significantly larger than the optimal area especially when $\alpha$ is 80%. As $\alpha$ increases, the difference between our area and the optimal area also increases. To guarantee reaching the confidence level, the neural networks tend to cover slightly more of the distribution. Since the distribution samples become sparser in the area far from the mean, even a small increase in coverage results in a significant increase in the area. Compared to the bounding box, while it may achieve a higher correctness rate due to its highly conservative nature, its area is significantly larger than that of the proposed method. Meanwhile, although the area of ILP is similar to that of NN, its correctness rate is significantly lower. Table IX shows the performance of the neural network on fanshape distribution. The neural network achieves a near-perfect correctness rate of close to 100%. Additionally, the area comparison follows a similar trend to that observed in the 2d binomial distribution. Table. IX shows the performance of the neural network on GMR trajectory predictor distribution. Although the correctness rate is not as high as in the previous tables, it still reaches a minimum of 93.1%. Meanwhile, the area remains superior to the Bounding Box, comparable to ILP, and only slightly larger than the optimal solution.
In summary, the proposed neural network demonstrates consistently high correctness rates across various distributions, ranging from at least 93.1% to nearly 100%, while maintaining a compact area that is only marginally larger than the optimal solution. This balance between correctness and efficiency underscores the utility of the neural network-generated convex sets. The result supports that the Advantage 1 of the proposed method, NN can generate useful PRS with only a few samples.
V-C Generating speed
In addition to comparing the accuracy, we compare the generating time required to generate useful results. Fig. 8 illustrates the average time cost for generating effective outcomes. The results clearly demonstrate that our method significantly outperforms the baseline approaches, achieving execution times that are an order of magnitude faster. The result supports that the Advantage 2 of the proposed method, NN can generate useful PRS very fast.
Figure 8: Generating speed comparison
VI Conclusion and Future
In this paper, we propose using self-supervised learning to train a transformer-based neural network to generate convex approximations of probabilistic reachable sets. Once trained, the neural network can robustly generate convex approximations that cover the desired percentage of points sampled from the distribution of the probabilistic reachable set only given a few samples from the distribution. The primary advantage of the proposed method is its extremely fast generation time while maintaining performance comparable to that of the baseline methods. Meanwhile, training is also very simple because training inputs can be generated very fast. Only a few experts are required for the labels. No physical information about the system such as disturbance is required.
Acknowledgment
The authors would like to acknowledge the support from the National Science Foundation under Grants CCF-2402689 and CMMI-2138612. Any opinions, findings, conclusions, or recommendations expressed in this paper are those of the authors and do not reflect the views of NSF.

### v2

@@
-Convex Approximation of Probabilistic Reachable Sets from Small Samples Using Self-supervised Neural Networks
+Learning Probabilistic Obstacle Spaces from Data-driven Uncertainty using Neural Networks
 Abstract
-Probabilistic Reachable Set (PRS) plays a crucial role in many fields of autonomous systems, yet efficiently generating PRS remains a significant challenge.
-This paper presents a learning approach to generating 2-dimensional PRS for states in a dynamic system.
-Traditional methods such as Hamilton-Jacobi reachability analysis, Monte Carlo, and Gaussian process classification face significant computational challenges or require detailed dynamics information, limiting their applicability in realistic situations.
-Existing data-driven methods may lack accuracy.
-To overcome these limitations, we propose leveraging neural networks, commonly used in imitation learning and computer vision, to imitate expert methods to generate PRS approximations.
-We trained the neural networks using a multi-label, self-supervised learning approach.
-We selected the fine-tuned convex approximation method as the expert to create expert PRS.
-Additionally, we continued sampling from the distribution to obtain a diverse array of sample sets.
-Given a small sample set, the trained neural networks can replicate the PRS approximation generated by the expert method, while the generation speed is much faster.
-Note to PractitionersReachability analysis is one of the popular methods for solving real-time motion planning problems for dynamic systems under uncertainty while ensuring safety.
-However, determining the reachable set is challenging.
-Traditional methods are computationally expensive and often require extensive system information.
-Meanwhile, neural networks have been shown to construct distributions effectively.
-Hence, we propose leveraging the capability of neural networks to model the distribution of reachable sets.
-To simplify the creation of training datasets, we incorporate self-supervised learning in training the neural network.
-The proposed method offers several advantages: it is easy to use, generates accurate results quickly, and requires less system information.
-However, as neural networks are statistical methods relying on data-driven insights, their effectiveness inherently depends on the availability of sufficient historical data for the system.
-Index Terms:
-Probabilistic reachable set, Convex approximation, Deep Learning
+Identifying the obstacle space is crucial for path planning.
+However, generating an accurate obstacle space remains a significant challenge due to various sources of uncertainty, including motion, behavior, and perception limitations.
+Even though an autonomous system can operate with an inaccurate obstacle space by being over-conservative and using redundant sensors, a more accurate obstacle space generator can reduce both path planning costs and hardware costs.
+Existing generation methods that generate high-quality output are all computationally expensive.
+Traditional methods, such as filtering, sensor fusion and data-driven estimators, face significant computational challenges or require large amounts of data, which limits their applicability in realistic scenarios.
+In this paper, we propose leveraging neural networks, commonly used in imitation learning, to mimic expert methods for modeling uncertainty and generating confidence regions for obstacle positions, which we refer to as the probabilistic obstacle space.
+The network is trained using a multi-label, supervised learning approach.
+We adopt a fine-tuned convex approximation method as the expert to construct training datasets.
+After training, given only a small number of samples, the neural network can accurately replicate the probabilistic obstacle space while achieving substantially faster generation speed.
+Moreover, the resulting obstacle space is convex, making it more convenient for subsequent path planning.
 I Introduction
-Autonomous system operations face various uncertainties and hazards [1].
-The reachable set approach is an important method in collision avoidance [2, 3], safety verification [4] and path planning [5] for Autonomous systems.
-However, finding the deterministic reachable sets with 100% confidence level is difficult [6].
-Meanwhile, the deterministic reachable sets may be too large and over-conservative to be applicable [7].
-Finding the Probabilistic Reachable Set (PRS) is more realistic, incorporating uncertainties and stochastic elements in the system dynamics or initial conditions [8].
-A brief visualization of PRS is shown in Fig.
-1, where PRS with different confidence levels are colored.
-Figure 1: Probabilistic Reachable set represents the set of possible states a dynamical system can reach from an initial state.
-For instance, a PRS with a 95% confidence level means that 95% of the possible system trajectories, considering uncertainties or noise, will remain within the set bounds
-I-A Review
-Finding PRS remains challenging.
-One method for determining reachable sets under bounded uncertainties is the Hamilton-Jacobi (HJ) reachability analysis.
-This approach formulates the problem as a partial differential equation that can be addressed using numerical techniques.
-Although this method is manageable in scenarios with fewer dimensions, its efficiency significantly diminishes as the dimensionality increases [9][10].
-Additionally, HJ reachability analysis requires detailed information about the dynamics or uncertainties.
-Meanwhile, strong assumptions are required, such as the absence of time correlations in parameter uncertainties.
-These limitations make it unsuitable for certain realistic situations [9].
-Fortunately, we can identify the reachable sets using a learning-based method given a large dataset from an unknown distribution.
-Generating a reachable set with machine learning/data-driven has been explored.
-A trained locally-weighted regression algorithm and nonlinear SVM classifiers can predict the reachability of the two-point boundary value problem (2PBVP) [11].
-Other data-driven methods such as the Gauss process [12], Monte Carlo [13], and stochastic disturbance [14] have also been developed to approximate reachable set.
-However, all of those methods have significant limitations in practice for three aspects: a) they all require a lot of samples to perform; b) they cannot provide fast and accurate probabilistic reachable set with adjustable confidence levels; c) their PRS bound is not convex and linear.
-Deep learning technology has seen significant advancements recently in computer vision [15] and nature language process [16].
-And those methods are further adopted in many different fields such as finance [17], cyber security [18], healthcare [19, 20] and robotics [21, 22, 23, 24].
-Deep learning-based methods have been effectively applied to solve obstacle avoidance [25] and path planning problems [26, 27], demonstrating their suitability for addressing challenges in autonomous control tasks.
-The Universal Approximation Theorem [28], guarantees that a feedforward neural network can approximate any nonlinear real-valued function.
-The advantages of neural networks include their ability to generate complex outputs from simple inputs, as seen in generative models.
-Previous research shows that neural networks can work when massive data is massively missing [29] and can determine uncertainty [30].
-Neural networks can rapidly extract the texture information of pavement from just a few images [31].
-Neural networks indeed offer significant advantages in terms of reduced computing time.
-Also, the parallel nature of neural networks can lead to significant reductions in computing time [32].
-Furthermore, Transformers [16] and self-supervised learning have revolutionized the field of artificial intelligence [33] and have become essential tools for developing more efficient, and versatile AI models when limited sample is available [34].
-Therefore, we propose designing a transformer-based neural network to generate PRS, and we further employ self-supervised learning techniques to train this network.
-I-B Contribution
-In this paper, we propose a method that uses self-supervised and multi-label training to train a transformer-based neural network for generating convex approximations of PRS from a small number of samples.
-The numerical experiments prove that the neural network can generate reliable convex approximations when the given samples come from a distribution that has been learned by the network.
-This effective method relies on just one assumption:
-  • Assumption 1: a large number of previous samples have been observed before the predicting task.
-This assumption is satisfied in many fields such as autonomous driving [35], flight landing prediction [36], and robotics [37].
-Our method offers three unique advantages that set it apart from other methods:
-  • Advantage 1: No information about the desired system is needed; only a few samples of the desired distribution are required for PRS generating.
-  • Advantage 2: The generation speed is extremely fast, making it suitable for online applications.
-  • Advantage 3: All the PRS bound generated are convex and linear.
-  • Advantage 4: Very few labeled data are needed for training, which simplifies the training process and makes large-scale training feasible.
-II Problem Modeling
-The main goal of the paper is to generate a convex approximation that can capture the PRS of an unknown static uncertainty whose data samples are gradually observed.
-In this section, we will introduce and discuss the concepts of reachable sets and PRS associated with a dynamic system with uncertain discrete-time characteristics  [38, 39].
-Let us consider a general discrete-time dynamic system expressed as
-$\boldsymbol{s}_{k+1}=\boldsymbol{f}\left(\boldsymbol{s}_{k},\boldsymbol{u}_{k},\boldsymbol{\theta}_{k},\boldsymbol{w}_{k}\right),$ (1)
-where, $\boldsymbol{s}_{k}\in\mathbb{R}^{2}$ denotes the state of the system (we only consider 2d dynamic systems in this paper, we use $s$ instead of $x$ like many other dynamic system research to distinguish x-dimension) at time step $k\in\{1,\ldots,T\}$, $\boldsymbol{u}_{k}\in\mathcal{U}_{k}$ denotes an uncertain control input, $\boldsymbol{\theta}_{k}\in\Theta$ denotes an uncertain parameter, $\boldsymbol{w}_{k}\in\mathbb{W}$ denotes an uncertain disturbance, and $s_{0}\in\mathcal{S}_{0}$ denotes the initial state.
-Diverging from most previous studies [9], this paper allows the sets $\mathcal{U}_{k}\subset\mathbb{R}^{m}$, $\Theta\subset\mathbb{R}^{p}$, $\mathbb{W}\subset\mathbb{R}^{q}$, and $\mathcal{S}_{0}\subset\mathbb{R}^{2}$ to be unbounded.
-Taking into account the uncertainties in $\boldsymbol{s}_{k}$, $\boldsymbol{u}_{k}$, $\boldsymbol{\theta}_{k}$, and $\boldsymbol{w}_{k}$, it follows that $\boldsymbol{x}_{k+1}$ becomes a random vector following an unknown probability distribution.
-Then at time $T$, the reachable set $\mathcal{S}_{T}$ of the dynamic system in Eq.
-1 is defined to be
-[…]
-$\boldsymbol{s}_{0}\in\mathcal{S}_{0},\boldsymbol{u}_{k-1}\in\mathcal{U}_{k-1},\boldsymbol{\theta}_{k-1}\in\Theta,\boldsymbol{w}_{k-1}\in\mathbb{W},$
-$k\in\{1,\ldots,T\}\},$
-where $\boldsymbol{s}_{0}$ denotes the initial state and $\mathcal{S}_{0}$ denotes the initial set.
-Observe that because the sets $\mathcal{U}_{k-1}$, $\Theta$, $\mathbb{W}$, and $\mathcal{S}_{0}$ can be unbounded, the resulting reachable set $\mathcal{S}_{T}$ could also be unbounded.
-This implies that determining a bounded set where the state is definitively contained might not be feasible.
-Therefore, our goal is to generate a bounded set where the probability of the state being within this set exceeds a certain confidence level, which is called a Probabilistic Reachable Set (PRS).
-At time $k$, a bounded set $\tilde{\mathcal{S}}_{k}$ can be qualified as a PRS of the dynamic system defined in Eq.
-1 at confidence level $\alpha$ if and only if
-$\operatorname{Pr}\left(\boldsymbol{s}_{k}\in\tilde{\mathcal{S}}_{k}\right)\geq\alpha.$ (3)
-It indicates that the probability of the system reaching the PRS at time $k$ exceeds $\alpha$.
-For example, if $\alpha=95\%$ and the system runs an infinite number of times, the system will be outside the PRS less than 5% of the time.
-When $\alpha=100\%$ and a PRS $\tilde{\mathcal{S}}_{k}$ does exist, then the reachable set $\mathcal{S}_{k}$ defined in Eq.
-2 belongs to the $\tilde{\mathcal{S}}_{k}$.
-The previous expert method [40] can find multiple $\tilde{\mathcal{S}}_{k}$ and all the founded $\tilde{\mathcal{S}}_{k}$ are convex and as tight as possible.
-We define the set $\tilde{\mathcal{S}}_{k}$ using the positions of the four vertices, represented as $\tilde{\mathcal{V}}_{k}\in\mathbb{R}^{4\times 2}$.
-The trained neural network should be able to imitate the expert and generate a similar PRS that has a confidence level close to $\alpha$, as shown in Eq.
-4,
-$\tilde{\mathcal{V}}_{k}=NN(w,\mathbf{s_{k}})$ (4)
-where $w$ is the trained weights of the neural network, $\mathbf{s_{k}}=\{s_{k}^{i}\}_{i=1}^{n}$ is a collection of samples from the real distribution of the reachable set with $s_{k}^{i}\sim\boldsymbol{s}_{k}$, and $n$ as the number of the samples.
-The desired $n$ should be smaller than 100.
-In the rest of the paper, we only consider the reachable set for a specific single time step, for example, $k=T$.
-So to simplify the notation, we use $\mathcal{S}$ to refer the actual reachable set, $\tilde{\mathcal{S}}$ to refer the PRS, $\tilde{\mathcal{V}}$ as the four vertices of the PRS, and $\boldsymbol{s}$ as the state of the system, $\mathbf{s}$ as the collection of samples from the reachable set at desired time point.
-III Dataset and Expert Method
-(a) Example of 2d-binominal distribution and sample expert convex PRS
-(b) Example of fan-shape distribution and sample expert convex PRS
-(c) Example of GMR trajectory predictor distribution and sample expert convex PRS
-Figure 2: Distribution and convex PRS demo
-III-A Target distribution
-In this paper, we evaluate the proposed method to determine if it can approximate three different probabilistic distributions, all of which are based on the Gaussian distribution, commonly used in aircraft trajectory modeling [41].
-Any sample from the target probabilistic distributions is 2-dimensional data, $\mathbf{s}\in R^{2}$.
-III-A1 2d-binomial distribution
-2d-binomial distribution [38] forms clusters around two different coordinate points.
-The sampling plot of the 2d-binominal distribution is shown in Fig.2(a).
-The sample process includes two steps, first is randomly selecting one of two center points, second is adding Gaussian noise to both the x- and y-coordinates of the chosen center point.
-Let $\boldsymbol{C}_{1}=(x_{1},y_{1})$ and $\boldsymbol{C}_{2}=(x_{2},y_{2})$ represent the two center points.
-Let $\mathbf{N}=\begin{pmatrix}N_{x}\\ N_{y}\end{pmatrix}$ be the Gaussian noise added to both coordinates, where $N_{x}\sim\mathcal{N}(0,\sigma^{2})$ and $N_{y}\sim\mathcal{N}(0,\sigma^{2})$ are independent Gaussian random variables with mean 0 and variance $\sigma^{2}$.
-The sample of the 2d-binomial distribution will be $\boldsymbol{s}=\boldsymbol{C}_{i}+\mathbf{N},i\in\{1,2\}$.
-The probability density function(PDF) of the 2d-binomial distribution is easy to calculate:
-$P(\boldsymbol{s})=\frac{1}{2}\cdot\mathcal{N}(\boldsymbol{s}|\boldsymbol{C}_{1},\sigma^{2}\mathbf{I})+\frac{1}{2}\cdot\mathcal{N}(\boldsymbol{s}|\boldsymbol{C}_{2},\sigma^{2}\mathbf{I}),$ (5)
-where, $\mathcal{N}(\boldsymbol{s}|\boldsymbol{C}_{i},\sigma^{2}\mathbf{I})$ represents the bivariate Gaussian distribution centered at $\boldsymbol{C}_{i}=(x_{i},y_{i})$ with covariance matrix $\sigma^{2}\mathbf{I}$ (identity matrix scaled by $\sigma^{2}$, meaning independent noise with the same variance for $x$ and $y$).
-III-A2 Fanshape distribution
-The fan-shape distribution[38] represents points sampled from a distribution of motion paths influenced by truncated normal distributions of speed and direction, capturing realistic variations in both parameters.
-The sampling plot of the fanshape distribution is shown in Fig.2(b).
-[…]
-The position $\mathbf{s}:=[x_{k},y_{k}]^{\top}$ of the flight at time $k$ is modeled by the following discrete-time dynamic system:
-$\begin{bmatrix}x_{k}\\ y_{k}\end{bmatrix}=\begin{bmatrix}x_{k-1}\\ y_{k-1}\end{bmatrix}+\begin{bmatrix}\cos\theta\\ \sin\theta\end{bmatrix}v\Delta t,$
-where $[x_{k-1},y_{k-1}]^{\top}$ is the position of the fight at the previous time step $k-1$.
-$[x_{k},y_{k}]^{\top}$ is the position of the flight at the time step $k$.
-$v$ is the speed and $\theta$ is heading angle.
-Due to the uncertainties in $v$ and $\theta$, the position $\mathbf{s}$ is a random vector that obeys a non-Gaussian distribution, which is referred to as the fanshaped distribution.
-The PDF of the position is difficult to derive.
-III-A3 GMR trajectory predictor
-GMR trajectory predictor can regress the possibility distribution of the future position of flight near terminal [42] with multiple Gaussian distributions given the past trajectory.
-The sampling plot of the GMR trajectory predictor distribution is shown in Fig.2(c).
-The regressed conditional possibility distribution of the future position of flight is defined by the following equation:
-$P(\boldsymbol{s}\mid\boldsymbol{s}^{past})=\sum_{k=1}^{K}\pi_{\boldsymbol{s}\mid\boldsymbol{s}^{past}_{k}}\mathcal{N}_{k}\left(\boldsymbol{s}\mid\boldsymbol{\mu}_{\boldsymbol{s}\mid\boldsymbol{s}^{past}_{k}},\boldsymbol{\Sigma}_{\boldsymbol{s}\mid\boldsymbol{s}^{past}_{k}}\right),$ (6)
-where where $\mathcal{N}_{k}\left(\boldsymbol{s}\mid\boldsymbol{\mu}_{\boldsymbol{s}\mid\boldsymbol{s}^{past}_{k}},\boldsymbol{\Sigma}_{\boldsymbol{s}\mid\boldsymbol{s}^{past}_{k}}\right)$ are Gaussian distributions with mean $\boldsymbol{\mu}_{\boldsymbol{s}\mid\boldsymbol{s}^{past}_{k}}$ and covariance $\boldsymbol{\Sigma}_{\boldsymbol{s}\mid\boldsymbol{s}^{past}_{k}}$, $K$ is the number of Gaussians, and, $\boldsymbol{s}^{past}_{k}$ is the input and $\boldsymbol{s}$ is the output, where $\boldsymbol{s}^{past}_{k}$ is the combined past trajectory and future trajectory guide of the ego agent.
-$\boldsymbol{s}$ is the trajectory prediction.
-$\pi_{k}\in[0,1]$ are weights that sum up to one.
-If $\pi_{k}$ is larger, the more likely the $\boldsymbol{s}$ belongs to $\mathcal{N}_{k}$.
-We use three-time step prediction as the desired distribution.
-III-B Expert method
-The expert method is proposed in [40].
-The expert method can generate the PRS of an uncertain dynamic system.
-This data-driven method builds on a kernel density estimator (KDE) that can generate the PRS, accelerated by a fast Fourier transform (FFT).
-These KDE values are then utilized within an Integer Linear Programming (ILP) framework to find a minimal parallelogram box that approximates the probabilistic geofence.
+Autonomous vehicles operations face various uncertainties and hazards [1].
+One of the most important tasks of any autonomous vehicle is to move without collisions, in other words, to remain within a safe space at all times [2].
+Safe space is also called safe region [3], obstacle-free space [4] in other works.
+The first step to find a safe space is usually to find the obstacle space.
+While many methods exist for generating a safe space under the assumption of a known obstacle space [5], this assumption is often unrealistic in practice, and only a few approaches [6, 7] extend to uncertain obstacles.
+Meanwhile, many previous path planning algorithms assume the obstacle space is polygonal [8, 9] due to the simplicity of collision checking and the compactness of the representation.
+The area of the obstacle space should be minimized, as a larger safe space typically results in smoother and shorter trajectories while also reducing the overall planning cost [4].
+Building on these insights, this paper proposes a novel method to construct a polygonal probabilistic obstacle space that is both compact and capable of representing uncertain obstacles.
+In practice, safety is ensured by declaring a collision whenever the distance between an agent and an obstacle falls below a threshold, even without physical contact [10].
+In this work, we focus on obstacles with relatively small volumes, such as flights or UAVs in airspace [11], where the obstacle space is determined primarily by position rather than precise geometry, a common simplification known as the point mass model.
+An obstacle is certain if a deterministic value represents its position, and uncertain if a probability distribution describes its position.
+The probabilistic obstacle space is then defined as the confidence region of this distribution.
+Uncertainty of an obstacle can be caused by many reasons.
+The first common reason is perception uncertainty.
+All of the position-locating sensors, including GPS [12], camera [13], Lidar [14], and radar [15], are subject to measurement noise.
+It is necessary to construct a probabilistic obstacle space from noisy measurements.
+Sensor Uncertainty Mitigation (SUM) methods [16] can compute bounds on the errors, but it has a very high computational cost and require diverse and representative data.
+Meanwhile, those over-conservative approaches make safe spaces very small, making it difficult for planning algorithms to find feasible solutions
+Behavioral uncertainty can also render an obstacle uncertain, since different intents lead to entirely different trajectories [17].
+Traditional data-driven methods [18, 19] can estimate intent distributions when sufficient historical data are available.
+However, such distributions are often highly complex and require many parameters to represent, which makes it impractical to directly translate them into obstacle spaces.
+Another common source of uncertainty is motion.
+Even obstacles following a fixed path can exhibit positional variability [6, 20].
+For instance, as shown in Fig.
+2, a flight may follow its planned trajectory yet still wander around it.
+Consequently, even with state-of-the-art trajectory predictors, the error between the predicted and actual positions can never be eliminated.
+The M-estimator of Tukey [21] has been used to estimate obstacle motion from single-camera images, but it is computationally expensive and requires careful parameter tuning.
+In this paper, we propose a supervised learning approach that employs multi-label training to train a transformer-based neural network for generating polygonal probabilistic obstacle spaces under uncertainty.
+Numerical experiments demonstrate that the network can reliably reproduce probabilistic obstacle spaces when the input samples are drawn from obstacle types it has learned.
+In addition, the proposed method achieves substantially faster generation times compared to baseline approaches.
+The method offers four key advantages:
+  • Advantage 1: The approach does not rely on dynamic models of uncertain obstacles.
+Instead, it can construct obstacle spaces from only a few noisy measurements or partial observations, reducing the sensing and modeling requirements.
+  • Advantage 2: All generated obstacle spaces are convex and linear, which simplifies their integration into path planning algorithms and ensures computational tractability.
+  • Advantage 3: Training requires only a small amount of labeled data, which lowers the burden of data collection and annotation and makes the approach practical for large-scale deployment.
+  • Advantage 4: The generation process is extremely fast, enabling real-time use in online planning and decision-making systems.
+II Problem Statement
+The main goal of this paper is to generate the probabilistic obstacle space of an uncertain obstacle from a limited data sample using a neural network:
+$o^{\alpha}=NN(w,\mathbf{s}),$ (1)
+where $o^{\alpha}$ is the probabilistic obstacle space at confidence level $\alpha\in(0,1)$, $w$ denotes the neural network parameters, and $\mathbf{s}={s^{i}}_{i=1}^{n}$ is the set of observed $n$ samples.
+Definition (Probabilistic Obstacle Space).
+Let $o^{true}$ be the true occupied region of an uncertain obstacle $o$, and let $\mathbb{P}$ denote the underlying probability measure that characterizes uncertainty in the obstacle’s position.
+A set $o^{\alpha}$ is called a probabilistic obstacle space at confidence level $\alpha$ if it satisfies
+$\mathbb{P}[a\in o^{true}\mid a\notin o^{\alpha}]<1-\alpha.$ (2)
+That is, if the agent $a$ avoids $o^{\alpha}$, then the probability of entering the true obstacle region $o^{true}$ is bounded above by $1-\alpha$.
+Since $o^{true}$ is generally unknown due to imperfect perception and prediction, $o^{\alpha}$ provides a tractable confidence-region approximation.
+Our definition generalizes several prior notions, including the $\epsilon$-shadow of obstacle spaces [6, 7], Gaussian-distributed obstacles [22], adaptive bounding boxes [23], and $\alpha$-probabilistic occupied sets [24].
+The neural network in Eq. (1) is thus trained to approximate the mapping
+$f:\mathcal{S}^{n}\to\mathcal{O},\quad f(\mathbf{s})\mapsto o^{\alpha},$ (3)
+where $\mathcal{S}^{n}$ is the sample space of $n$ measurements and $\mathcal{O}$ is the family of convex polygonal sets in $\mathbb{R}^{d}$.
+We illustrate three representative examples of probabilistic obstacle spaces and their associated data samples.
+Figure 1: Motion uncertainty: even with a planned trajectory, the actual obstacle position is affected by disturbances and uncertainties.
+Motion uncertainty: Consider the scenario of a moving object following a fixed planned trajectory.
+However, many factors, such as environmental disturbances, control uncertainties, and human factors, can introduce uncertainty in the actual trajectory.
+Therefore, as shown in Fig.
+1, a single planned trajectory may correspond to infinitely many possible realizations of motion, all of which the ego agent must avoid.
+To capture this variability, we construct a probabilistic obstacle space that encompasses the possible deviations.
+Formally, we define the sample set $\mathbf{s}={s^{i}}_{i=1}^{n}$ as the trajectories of $n$ flights following the same planned path.
+Figure 2: Behavior uncertainty: flights taking off from the same runway with a different destination will have a different route
+Behavior uncertainty: Another important source of uncertainty arises from the intent of other agents.
+Consider an ego UAV operating in terminal airspace near a busy airport.
+Although aircraft follow air traffic control clearances and standard procedures, the UAV does not know in advance which clearance each departure will receive.
+As shown in Fig.
+2, multiple aircraft may depart from the same runway within a narrow time window, and each departure may be cleared for a different standard instrument departure depending on destination, weather, and traffic flow.
+From the UAV’s perspective, the resulting obstacle space is therefore governed by a probability distribution over several candidate routes rather than a single deterministic path.
+Because such intent distributions are complex and rarely available in closed form, we approximate them by constructing the probabilistic obstacle space from $\mathbf{s}$, defined as the trajectories of $n$ flights that have previously departed from the same runway.
+Sensor uncertainty: Fig.
+3 illustrates two adjacent circular obstacles whose positions are estimated using noisy sensor data.
+Because the measurements are corrupted by noise, the exact obstacle positions cannot be determined without calibration or advanced signal processing techniques.
+Furthermore, a single measurement outcome may be consistent with multiple possible true obstacle positions.
+To account for this, we define $\mathbf{s}$ as the set of $n$ noisy sensor measurements, from which a probabilistic obstacle space is constructed.
+Figure 3: Sensor uncertainty: the position of two adjacent obstacles whose positions are estimated using noisy sensor measurements is uncertain
+Since the obstacle space is inherently probabilistic, its exact boundary may be irregular or complex.
+For path planning, however, a tractable geometric representation is preferred.
+In particular, polygonal approximations are widely used because they provide a convex, linear structure that can be efficiently integrated into optimization-based planners.
+To this end, instead of directly generating the full probabilistic obstacle space, we represent it using a convex polygon defined by its vertices.
+Specifically, the neural network outputs four vertices that characterize the polygonal approximation of the probabilistic obstacle space:
+$\tilde{\mathcal{V}}=NN(w,\mathbf{s}),$ (4)
+where $w$ denotes the trained parameters of the neural network and $\tilde{\mathcal{V}}$ is the predicted set of polygon vertices.
+III Method
+III-A Supervised learning and expert method
+Figure 4: Supervised learning: we iteratively pick a few samples from $o^{true}$ as inputs to the neural network and task the network with generating the corresponding probabilistic obstacle space.
+Previous research [25] has proven that the transformer can process the visible (unmasked) parts of the data, reconstructing the entire input from the encoded, incomplete data (masked data).
+Therefore, if we only have a few samples from a distribution, the transformer may be able to reconstruct the entire distribution.
+To train the neural network, we first require expert solutions generated by an expert method to serve as supervision during the training process.
+In this paper, we employ the expert method, as proposed in our previous work [26].
+The expert method can generate the probabilistic obstacle space of an uncertain object.
+This data-driven method builds on a kernel density estimator (KDE) that can generate the probabilistic obstacle space, accelerated by a fast Fourier transform (FFT).
+These KDE values are then utilized within an Integer Linear Programming (ILP) framework to find a minimal parallelogram box that approximates the probabilistic obstacle space.
 This ILP solution effectively encapsulates the area where the state is expected to stay with high confidence.
@@
 The corresponding $z_{ij}$ entry is set to one within these boundaries, yielding the matrix as an efficient approximation of the ILP solution.
-The expert convex generated is shown in Fig.
-2.
-The black points represent the distribution and purple convex covers a desired percent of the distribution.
-III-C Training dataset generation
-Theoretically, there are unlimited correct convex approximations for PRS given a probability distribution.
-At the same time, the expert method introduced in the previous section can generate a very accurate convex approximation if enough samples are given.
-Therefore, for each distribution $P$ and each disired confidence level $\alpha$, we can generate a set of correct convex approximation $\mathbf{\tilde{\mathbf{S}}}_{p,\alpha}=\{\tilde{\mathcal{S}}_{P,\alpha}^{i}\}_{i=1}^{n}$.
-In this paper, the correct convex approximation has 1% error toleration.
-Therefore, we consider the approximation is correct if $\alpha+0.01\geq\operatorname{Pr}\left(\boldsymbol{s}\in\tilde{\mathcal{S}}_{\alpha}\right)\geq\alpha$.
-Each correct convex approximation $\tilde{\mathcal{S}}_{P,\alpha}$ is defined by four vertices $\tilde{\mathcal{V}}_{P,\alpha}$.
-Then the set of vertices of the correct convex approximation is denoted as $\mathbf{\tilde{\mathbf{V}}}_{p,\alpha}=\{\tilde{\mathcal{V}}_{P,\alpha}^{i}\}_{i=1}^{n}$.
-The usage of the generated correct convex approximation will be introduced in the method section below.
-IV Method
-IV-A Self-supervised learning
-Figure 3: Masked distribution: we only show a few samples to the neural networks, and ask neural networks to generate the corresponding PRS.
-Previous research [33] has proven that the transformer can process the visible (unmasked) parts of the data, reconstructing the entire input from the encoded, incomplete data (masked data).
-Therefore, if we only have a few samples from the distribution, the transformer may be able to reconstruct the entire distribution.
-As the previous section III, we have the vertices of expert convex approximation $\mathbf{\tilde{\mathbf{V}}}_{p,\alpha}$.
-Fig.3 shows the examples of how we mask the distribution.
-During the training, we randomly mask most of the sample points, only leaving a few samples visible to the transformer, and ask the transformer to generate the corresponding PRS.
-With this self-supervised learning strategy, we are able to show many sets of visible samples of one distribution to the neural networks.
-When we show the stranger samples to the trained neural networks, the trained neural network is able to identify the distribution by recognizing and relating the visible samples to the samples the neural networks have seen during the training process.
-IV-B Multi-label learning and loss function
-Figure 4: Example of multi-label learning
-As introduced in section II, one distribution can have multiple correct convex approximations.
-Therefore, we can use multi-label learning instead of single-label learning.
+After obtaining the expert polygon vertices $\tilde{\mathcal{V}}$, we prepare the input data for network training.
+Fig.
+4 illustrates how different subsets of samples are presented to the neural network during training.
+At each training step, a small random subset of samples is selected, and the transformer is trained to generate the corresponding polygon vertices.
+Through this supervised learning process, the network is repeatedly exposed to multiple partial views of the same obstacle, learning to associate incomplete sample sets with the full probabilistic obstacle space.
+At inference time, when only a limited number of samples are available, the trained network can infer the underlying distribution of the obstacle space by leveraging the patterns it has learned from these partial-to-full mappings.
+III-B Multi-label learning and loss function
+Figure 5: Multi-label learning: during training, the network is guided to produce outputs that align with the closest expert solution.
+Recall the non-uniqueness of confidence regions [27]: for any given distribution, there exist infinitely many subsets that capture the same probability mass.
+Consequently, for each uncertain obstacle, there are infinitely many valid probabilistic obstacle spaces.
+This motivates the use of multi-label learning rather than single-label learning.
 Multi-label learning not only reduces computational overhead but also allows the model to share learned features across multiple labels, potentially leading to better generalization.
 Fig.
-4 shows an example of multi-label learning.
-The neural network generated an output based on the current weights of the neural network.
-Then the losses between each expert and the output are calculated and the smallest loss will be used to train the neural network.
-For the real training, we will have more expert PRS samples.
-Like many regression machine learning research studies, we use the MSE loss function as shown in Eq.
-7.
-MSE loss can help the neural network minimize the average error between the expert and the output.
-$\text{Loss}=\min_{j\in[1,10]}{\frac{1}{8}\sum_{i=1}^{8}(\hat{y}_{ij}-y_{i})^{2}}\quad\text{where}\quad\hat{y}_{ij}\in\mathbf{\tilde{\mathbf{V}}}_{p,\alpha}$ (7)
+5 illustrates this process.
+Given the current network parameters, the neural network produces an output polygon.
+The loss is then computed between this output and each available expert probabilistic obstacle space.
+The smallest loss among them is used to update the network weights.
+In practice, multiple expert probabilistic obstacle spaces are provided for each obstacle during training.
+Similar to many regression-based approaches, we employ the mean squared error (MSE) loss, as shown in Eq.
+5, which minimizes the average discrepancy between the network output and the expert labels.
+$\text{Loss}=\min_{j\in[1,10]}{\frac{1}{8}\sum_{i=1}^{8}(\hat{y}_{ij}-y_{i})^{2}}\quad\text{where}\quad\hat{y}_{ij}\in\mathbf{\tilde{\mathcal{V}}}$ (5)
 where:
@@
   • $\hat{y}_{ij}$ is the $j$-th expert vertices for the $i$-th element.
-IV-C Neural networks architecture
+III-C Neural networks architecture
 The overall architecture of the neural network we used is shown in Fig.
-5, which is designed based on our previous work [43].
+6, which is designed based on our previous work [28].
 The input to the neural network consists of sampled points and the desired confidence level.
@@
 In the transformer, the channel representing the confidence level interacts with all the point channels.
-Finally, we send the confidence level channel to an output layer, resulting in four vertices of the convex hull.
+Finally, we send the confidence level channel to an output layer, resulting in four vertices of the convex polygon.
 Figure 6: Neural network architecture
 Figure 7: Concatenated tensor
-IV-D Input normalization, embedding, and concatenate
-Because sample points are drawn from different distributions, they have very different means and variances, which can cause challenges for the training process of neural networks.
+III-D Input normalization, embedding, and concatenate
+Because sample points are drawn from different tasks, they have very different means and variances, which can cause challenges for the training process of neural networks.
 Normalizing those points before feeding them into the neural network is very crucial for several reasons.
-First, if certain inputs have larger values and higher variance than others, they may dominate the training process.
+First, if some inputs have larger values and higher variance than others, they may dominate the training process.
 For example, if one target (or feature) has a very large value compared to others, its gradient will be disproportionately large.
@@
 The concatenated tensor then looks like the Fig.
-6.
-IV-E Transformer
-For this work, we only use the transformer encoder [16] since we output the vertex at once instead of one by one.
-The encoder is composed of a stack of 6 identical layers, each layer has two sub-layers, a multi-head self-attention mechanism, and the feedforward network.
+7.
+III-E Transformer
+For this work, we only use the transformer encoder [30] because our task requires predicting all polygon vertices at once, instead of generating them sequentially.
+The encoder is composed of a stack of 6 identical layers, each layer has two sub-layers, a multi-head self-attention mechanism, and a feedforward network.
 We also kept the residual connection and layer normalization.
@@
 In contrast, each point interacts with all the other points at the same time in self-attention.
-IV-F Output
-In the previous masked autoencoder research.
-Mask tokens are introduced after the encoder, so the decoder will know which missed position is required to be generated.
-However, unlike reconstructing assignments, we do not have mask tokens.
-We do not know the visible samples are from which part.
-In fact, unlike CV and NLP problems, the sequence does not matter in our problem.
-For example, ”The dog chased the cat” and ”The cat chased the dog” are two completely different sentences, even though they have exactly the same words.
-In contrast, in our problem, the point set [(1, 1), (2, 2)], and [(2,2), (1, 1)] have the same distribution.
-Therefore, no additional global context is required.
-Furthermore, the goal of the proposed neural network is to generate a convex bound instead of reconstructing the whole distribution.
-All in all, a simple output layer is enough to generate the convex bound with the features captured from the visible samples.
+III-F Output
+In prior masked autoencoder research, mask tokens are introduced after the encoder so that the decoder can identify which missing positions need to be reconstructed.
+Our setting is fundamentally different.
+Unlike reconstruction tasks, we do not use mask tokens, since the visible samples do not carry positional information about which part of the obstacle space they come from.
+Moreover, sequence order is irrelevant in our problem.
+In natural language processing (NLP), for example, the sentences ‘The dog chased the cat’ and ‘The cat chased the dog’ convey completely different meanings despite containing the same words.
+In contrast, for our problem, the point sets (1,1),(2,2){(1,1),(2,2)} and (2,2),(1,1){(2,2),(1,1)} represent the same distribution.
+Consequently, no additional global context is required.
+Furthermore, the objective of our neural network is not to reconstruct the full underlying distribution but to generate a convex bounding region.
+Thus, a simple output layer suffices to produce the convex bound from the features extracted by the encoder.
 We use a multilayered output like ViT [31] instead of a single linear layer like the original transformer.
-The multi-layer output can further process the features extracted from the points channels and refine the final classification.
-We only used the transformed confidence level channel to output.
+The multi-layer output can further process the features extracted from the point channels and refine the final classification.
+We only used the transformed confidence level channel for output.
 The transformer allows the confidence level channel to attend to all other channels (embedded points), gathering contextual information from all the points.
 Finally, we need to denormalize the output, since the points we want to bound are normalized.
-IV-G Parameters
-For the transformer, embedding dimension $N_{emd}$ is 864.
+III-G Parameters
+For the transformer, the embedding dimension $N_{emd}$ is 864.
 We use 6 attention heads, so each head processes 144 embedding dimensions of each channel.
@@
 Additionally, a dropout rate of 0.2 is applied to prevent overfitting by randomly omitting certain neurons during training.
-V Result
-(a) 2d-binominal distribution
-(b) Fanshape distribution
-(c) GMR trajectory predictor distribution
-Figure 7: Example outputs if three methods take the same visible points, visible points are enlarged as yellow points for greater visibility.
-V-A Experiment setting
-We compare the proposed method with the ILP Heuristic (ILP), which is the expert method, and the bounding box (BB) method [45][46], a popular approach to finding a convex approximation of the PRS.
-Both baseline methods are KDE-based.
-Fig.
-7 are examples of point distribution and convex hulls generated by different methods given the same visible points.
-All the experiences are run in Intel(R) Core(TM) i9-12900KF and NVIDIA GeForce RTX 3090.
-V-A1 Training dataset
-In each training iteration, we randomly assign ten sets of expert vertices as the target and randomly pick 100 samples (300 for the GMR trajectory predictor) from the desired distribution as the neural network’s input.
-To accelerate the training process, we prepare 100 sets of expert vertices for each desired confidence level and prepare 320,000 sets of samples before training begins.
-To make the required expert vertices, it will cost experts about 24 hours for all distributions.
-In addition to expert vertices generating, we also need to generate training input samples before training.
-The sample generation time varies across different distributions.
-For the binomial distribution, it takes only about 2 minutes to generate 320,000 sets of samples.
-For the fanshape distribution, the process takes approximately 40 minutes to produce the same number of sample sets.
-In the case of the GMR trajectory predictor, it requires around 30 hours to generate 320,000 sets of samples.
-Generally, we can prepare the training dataset for a known distribution within a few days.
-It supports that the Advantage 4 of the proposed method, the training dataset is easy to make.
-Additionally, the proposed method exhibits Advantage 3: all PRS are guaranteed to be convex and linear, as the neural network is specifically designed to generate only the vertices of convex sets.
-V-A2 PRS generating
-We tested the trained neural networks by presenting them with 5,000 random tasks (1000 for each $\alpha$) in each distribution.
-Each task contains 100 (300 for GMR) points of the distribution and has never been seen by the neural network.
-The neural network will generate the corresponding PRS for those points.
-At the same time, we also generate PRS with BB and ILP given the same points as the baseline.
-However, since BB and ILP are not learning-based and do not require a training process, comparing them directly may not be entirely fair.
-To adjust for this, we also compared the results using PRS generated by BB and ILP with 20 times more samples to thoroughly evaluate our proposed method.
-Besides the baseline method, we also run the MINLP Optimal Algorithm [47] to determine the smallest area of a convex satisfies the coverage requirement.
-V-A3 PRS evaluating
-To evaluate a PRS, we first need to determine the coverage.
-To represent the real distribution, we sampled 100,000 points from the desired distribution.
-The resulting coverage was determined by counting the percentage of these sampled points that fell within the generated PRS.
-Mean Squared Error(MSE) is the metric we used to evaluate the accuracy with which the PRS approximates the desired confidence level:
+IV Result
+(a) motion uncertainty
+(b) behavior uncertainty
+(c) sensor uncertainty
+Figure 8: Black points are sampled from the real obstacle distribution.
+Colorful polygons are example outputs when the three methods take the same yellow visible points.
+Yellow points are enlarged for visibility.
+IV-A Experiment setting
+To evaluate the effectiveness of the proposed method, we conduct experiments on three types of uncertain obstacles: flying obstacles subject to motion uncertainty, obstacles with uncertain destinations, and obstacles whose positions are estimated using noisy sensor measurements.
+We feed the neural network with samples drawn from the true obstacle distribution and desired confidence level $\alpha$, then evaluate whether the generated probabilistic obstacle space achieves the desired coverage, i.e., whether it bounds the specified proportion of the distribution.
+Obstacles of the same type share the same underlying distribution but may exhibit different geometric shapes.
+IV-A1 Training
+We trained a neural network that can work with varying values of $\alpha$, where $\alpha$ can be 80%, 85%, 90%, 95% and 99%.
+To accelerate the training process, we prepare 100 sets of expert vertices for each desired confidence level and prepare 320,000 sets of 100 samples before training began.
+In each training iteration, we randomly select one set of samples and one of $\alpha$ as the neural network’s input, then randomly assign ten sets of corresponding expert vertices as the target.
+We can prepare the training dataset for a known uncertain obstacle within a few hours.
+It supports the Advantage 3 of the proposed method, the training dataset is easy to make.
+IV-A2 Interface
+We tested the trained neural networks by presenting them with 5,000 random tasks (1000 for each $\alpha$) from each type of uncertain obstacle.
+Each task contains 100 samples randomly generated from the real distribution of the uncertain obstacle and has never been seen by the neural network.
+The neural network generates the corresponding $o^{\alpha}$ for those samples.
+IV-A3 evaluating
+To evaluate the $o^{\alpha}$ generated by the neural network, we first need to determine the coverage.
+We approximate the true obstacle distribution using 100,000 samples, as the underlying distributions are typically empirical.
+The resulting coverage was computed as the percentage of samples that fell within the generated $o^{\alpha}$.
+We use mean squared error (MSE) to quantify the accuracy of the $o^{\alpha}$:
 $\text{MSE}=\frac{1}{n}\sum_{i=1}^{n}\left(\alpha_{i}-\hat{\alpha}_{i}\right)^{2}.$ (7)
-$\alpha$ is the coverage of the generated PRS and $\hat{\alpha}_{i}$ is the desired coverage.
+$\alpha_{i}$ is the real coverage of the generated $o^{\alpha}$ and $\hat{\alpha}_{i}$ is the desired confidence level (CL).
 n is 1000 because we test 1000 times for each $\alpha$.
 We also recorded the minimum and maximum coverage values to evaluate performance stability.
-While a low MSE indicates that the PRS is accurate, it does not guarantee correctness.
-Therefore, it is essential to compare the percentage of instances where the PRS exceeds the desired coverage.
-We called it the correctness rate.
-V-A4 Generating time
-Generating time is a crucial property in deciding whether a method can be used in online applications. Therefore, we compare the time to generate one correct PRS by each method.
-V-B Compare coverage accuracy, correctness, and area to baseline
-2d-binomial NN(our method) BB(100) ILP(100)
-confidence level MSE(range) MSE(range) MSE(range)
+While a low MSE indicates that the probabilistic obstacle space is accurate, it does not guarantee correctness.
+Therefore, it is essential to compare the percentage of instances where $\alpha_{i}$ exceeds $\hat{\alpha}_{i}$.
+We called it the correctness rate:
+$\text{Correctness Rate}=\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}\!\left[\alpha_{i}>\hat{\alpha}_{i}\right],$ (8)
+where $1[\bullet]$ is the indicator function(1 if the condition is true, 0 otherwise).
+IV-A4 baseline
+We compare the proposed method with the ILP Heuristic (ILP), which is the expert method, and the bounding box (BB) method [32][33], a popular approach to finding a polygonal approximation of the obstacle space.
+Both baseline methods are KDE-based.
+We generate $o^{\alpha}$ with BB and ILP for the same obstacle as the baseline.
+Fig.
+8 are examples of obstacle distribution and $o^{\alpha}$ generated by different methods given the same samples.
+However, since BB and ILP are not learning-based methods and do not require a training process, a direct comparison may not be entirely fair.
+To address this, we also provide the baseline methods with 20 times more samples to enable a more thorough evaluation of our approach.
+In addition to these baselines, we implement the MINLP Optimal Algorithm [34] to compute the smallest convex region that satisfies the coverage requirement.
+All the experiences are run in Intel(R) Core(TM) i9-12900KF and NVIDIA GeForce RTX 3090.
+TABLE I: MSE and range on motion uncertainty distribution compared to baseline with a small number of samples
+NN(our method) BB(100) ILP(100)
+CL MSE(range) MSE(range) MSE(range)
+80% 0.32%(78.95% - 80.44%) 11.98%(80.28% - 99.80%) 2.83%(75.04% - 91.87%)
+85% 0.25%(84.44% - 86.21%) 9.88%(85.39% - 99.91%) 2.95%(75.24% - 94.01%)
+90% 0.23%(89.40% - 90.91%) 7.36%(89.60% - 99.99%) 2.69%(75.63% - 98.43%)
+95% 0.49%(95.22% - 96.02%) 4.32%(92.60% - 100%) 2.20%(85.67% - 99.88%)
+99% 0.98%(99.96% - 99.99%) 1.33%(98.03% - 100%) 1.12%(97.86% - 100%)
+TABLE II: MSE and range on motion uncertainty compare to baseline with a large number of sample
+NN(our method) BB(2000) ILP(2000)
+CL MSE(range) MSE(range) MSE(range)
+80% 0.32%(78.95% - 80.44%) 12.73%(88.66% - 96.29%) 2.01%(75.36% - 84.46%)
+85% 0.25%(84.44% - 86.21%) 10.98%(91.93% - 98.82%) 2.31%(81.43% - 90.18%)
+90% 0.23%(89.40% - 90.91%) 8.63%(95.93% - 99.75%) 1.90%(87.40% - 94.16% )
+95% 0.49%(95.22% - 96.02%) 4.81%(98.59% - 100%) 2.72%(94.35% - 99.06%)
+99% 0.98%(99.96% - 99.99%) 0.99%(99.96% - 100%) 0.98%(99.19% - 100%)
+TABLE III: correctness rate/Probabilistic obstacle space area for each desired confidence level on motion uncertainty
+Correctness rate
+CL NN BB(100) BB(2000) ILP(100) ILP(2000) optimal(10000)
+80% 98.7% 100% 100% 57.7% 79.8% 100%
+85% 99.8% 100% 100% 52.4% 64.4% 100%
+90% 99.7% 99.8% 100% 52.8% 82.3% 100%
+95% 100% 99.4% 100% 74.8% 98.9% 100%
+99% 100% 61% 100% 59.7% 100% 100%
+CL Probabilistic obstacle space area
+80% 1415.07 2342.52 2309.52 1783.52 1514.04 1172.71
+85% 1762.30 2790.63 2815.55 2138.71 1902.27 1579.52
+90% 2229.01 3367.96 3470.37 2652.60 2365.27 2107.79
+95% 3530.87 4525.46 4576.97 3382.33 3374.61 2742.49
+99% 4833.31 5749.98 6127.58 4967.29 4820.67 3711.22
+TABLE IV: MSE and range on behavior uncertainty compare to baseline with a small number of sample
+NN(our method) BB(300) ILP(300)
+CL MSE(range) MSE(range) MSE(range)
+80% 1.79%(79.54% - 84.27%) 11.31%(84.04% - 96.41%) 2.74%(75.00% - 88.70%)
+85% 1.76%(82.34% - 88.25%) 9.56%(88.09% - 98.02%) 2.62%(75.31% - 90.94%)
+90% 1.13%(88.31% - 91.8%) 7.37%(93.11% - 99.40%) 1.76%(81.48% - 96.56%)
+95% 0.85%(94.14\$ - 97.83%) 4.09%(95.82% - 99.95%) 1.11%(90.61% - 98.80%)
+99% 0.58%(99.18% - 99.94%) 0.80%(98.42% - 100%) 0.40%(97.03% - 99.95%)
+TABLE V: MSE and range on behavior uncertainty compare to baseline with a large number of sample
+NN(our method) BB(6000) ILP(6000)
+CL MSE(range) MSE(range) MSE(range)
+80% 1.79%(79.54% - 84.27%) 10.47%(85.60% - 93.33%) 2.40%(75.00% - 81.90%)
+85% 1.76%(82.34% - 88.25%) 9.45%(90.69% - 96.35%) 3.01%(85.60% - 93.33%)
+90% 1.13%(88.31% - 91.8%) 7.44%(95.89% - 98.67%) 1.37%(85.62% - 94.55%)
+95% 0.85%(94.14\$ - 97.83%) 4.22%(97.98% - 99.67%) 0.74%(92.35% - 97.50%)
+99% 0.58%(99.18% - 99.94%) 0.91%(99.72% - 99.97%) 0.25%(98.28% - 99.61%)
+TABLE VI: correctness rate/Probabilistic obstacle space area for each desired confidence level on behavior uncertainty
+Correctness rate
+CL NN BB(300) BB(6000) ILP(300) ILP(6000) optimal(30000)
+80% 93.1% 100% 100% 10.8% 3.9% 100%
+85% 95.2% 100% 100% 27.9% 16.2% 100%
+90% 99.8% 100% 100% 69.2% 55.8% 100%
+95% 95.6% 100% 100% 68.2% 65.4% 100%
+99% 100% 99.7% 100% 56.5% 75.8% 100%
+CL Probabilistic obstacle space area
+80% 2.75E-3 3.64E-3 3.46E-3 2.59E-3 2.39E-3 2.15E-3
+85% 3.44E-3 4.41E-3 4.35E-3 2.99E3 2.89E-3 2.39E-3
+90% 4.41E-3 7.37E-3 5.81E-3 3.83E-3 3.67E-3 2.99E-3
+95% 5.69E-3 7.89E-3 7.98E-3 5.03E-3 4.87E-3 3.95E-3
+99% 9.21E-3 11.19E-3 11.16E-3 8.01E-3 7.63E-3 6.46E-3
+TABLE VII: MSE and range on sensor uncertainty distribution compare to baseline with a small number of sample
+NN(our method) BB(100) ILP(100)
+CL MSE(range) MSE(range) MSE(range)
 80% 0.42%(79.12% - 82.05%) 11.04%(87.56% - 99.42%) 3.66%(72.00% - 93.54%)
@@
 99% 0.87%(97.66% - 99.94%) 1.87%(97.09% - 99.99%) 0.80%(98.97% - 99.99%)
-TABLE I: MSE and range on 2d-binomial distribution distribution compare to baseline with a small number of sample
-fanshape NN(our method) BB(100) ILP(100)
-confidence level MSE(range) MSE(range) MSE(range)
-80% 0.32%(78.95% - 80.44%) 11.98%(80.28% - 99.80%) 2.83%(75.04% - 91.87%)
-85% 0.25%(84.44% - 86.21%) 9.88%(85.39% - 99.91%) 2.95%(75.24% - 94.01%)
-90% 0.23%(89.40% - 90.91%) 7.36%(89.60% - 99.99%) 2.69%(75.63% - 98.43%)
-95% 0.49%(95.22% - 96.02%) 4.32%(92.60% - 100%) 2.20%(85.67% - 99.88%)
-99% 0.98%(99.96% - 99.99%) 1.33%(98.03% - 100%) 1.12%(97.86% - 100%)
-TABLE II: MSE and range on fanshape distribution distribution compare to baseline with a small number of sample
-GMR trajectory predictor NN(our method) BB(300) ILP(300)
-confidence level MSE(range) MSE(range) MSE(range)
-80% 1.79%(79.54% - 84.27%) 11.31%(84.04% - 96.41%) 2.74%(75.00% - 88.70%)
-85% 1.76%(82.34% - 88.25%) 9.56%(88.09% - 98.02%) 2.62%(75.31% - 90.94%)
-90% 1.13%(88.31% - 91.8%) 7.37%(93.11% - 99.40%) 1.76%(81.48% - 96.56%)
-95% 0.85%(94.14\$ - 97.83%) 4.09%(95.82% - 99.95%) 1.11%(90.61% - 98.80%)
-99% 0.58%(99.18% - 99.94%) 0.80%(98.42% - 100%) 0.40%(97.03% - 99.95%)
-TABLE III: MSE and range on GMR trajectory predictor distribution compare to baseline with a small number of sample
-2d-binomial NN(our method) BB(2000) ILP(2000)
-confidence level MSE(range) MSE(range) MSE(range)
+TABLE VIII: MSE and range on sensor uncertainty distribution compare to baseline with a large number of sample
+NN(our method) BB(2000) ILP(2000)
+CL MSE(range) MSE(range) MSE(range)
 80% 0.42%(79.12% - 82.05%) 16.26%(93.24% - 98.43%) 4.12%(72.01% - 86.43%)
@@
 99% 0.87%(97.66% - 99.94%) 0.93%(99.73% - 99.99%) 0.86%(99.49% - 99.98%)
-TABLE IV: MSE and range on 2d-binomial distribution distribution compare to baseline with a large number of sample
-fanshape NN(our method) BB(2000) ILP(2000)
-confidence level MSE(range) MSE(range) MSE(range)
-80% 0.32%(78.95% - 80.44%) 12.73%(88.66% - 96.29%) 2.01%(75.36% - 84.46%)
-85% 0.25%(84.44% - 86.21%) 10.98%(91.93% - 98.82%) 2.31%(81.43% - 90.18%)
-90% 0.23%(89.40% - 90.91%) 8.63%(95.93% - 99.75%) 1.90%(87.40% - 94.16% )
-95% 0.49%(95.22% - 96.02%) 4.81%(98.59% - 100%) 2.72%(94.35% - 99.06%)
-99% 0.98%(99.96% - 99.99%) 0.99%(99.96% - 100%) 0.98%(99.19% - 100%)
-TABLE V: MSE and range on fanshape distribution distribution compare to baseline with a large number of sample
-GMR trajectory predictor NN(our method) BB(6000) ILP(6000)
-confidence level MSE(range) MSE(range) MSE(range)
-80% 1.79%(79.54% - 84.27%) 10.47%(85.60% - 93.33%) 2.40%(75.00% - 81.90%)
-85% 1.76%(82.34% - 88.25%) 9.45%(90.69% - 96.35%) 3.01%(85.60% - 93.33%)
-90% 1.13%(88.31% - 91.8%) 7.44%(95.89% - 98.67%) 1.37%(85.62% - 94.55%)
-95% 0.85%(94.14\$ - 97.83%) 4.22%(97.98% - 99.67%) 0.74%(92.35% - 97.50%)
-99% 0.58%(99.18% - 99.94%) 0.91%(99.72% - 99.97%) 0.25%(98.28% - 99.61%)
-TABLE VI: MSE and range on GMR trajectory predictor distribution compare to baseline with a large number of sample
-2d-binomial correctness rate
-confidence level NN BB(100) BB(2000) ILP(100) ILP(2000) optimal(10000)
+TABLE IX: correctness rate/Probabilistic obstacle space area for each desired confidence level on sensor uncertainty
+Correctness rate
+CL NN BB(100) BB(2000) ILP(100) ILP(2000) optimal(10000)
 80% 99.7% 100% 100% 31.4% 14.7% 100%
@@
 99% 98.7% 95.2 100% 99.9% 100% 100%
-confidence level PRS area
+CL Probabilistic obstacle space area
 80% 35.22 45.65 46.14 36.06 33.95 33.01
@@
 99% 99.03 101.49 106.67 95.28 96.10 73.96
-TABLE VII: correctness rate/PRS area for each desired confidence level on 2d-binomial distribution
-fanshape correctness rate
-confidence level NN BB(100) BB(2000) ILP(100) ILP(2000) optimal(10000)
-80% 98.7% 100% 100% 57.7% 79.8% 100%
-85% 99.8% 100% 100% 52.4% 64.4% 100%
-90% 99.7% 99.8% 100% 52.8% 82.3% 100%
-95% 100% 99.4% 100% 74.8% 98.9% 100%
-99% 100% 61% 100% 59.7% 100% 100%
-confidence level PRS area
-80% 1415.07 2342.52 2309.52 1783.52 1514.04 1172.71
-85% 1762.30 2790.63 2815.55 2138.71 1902.27 1579.52
-90% 2229.01 3367.96 3470.37 2652.60 2365.27 2107.79
-95% 3530.87 4525.46 4576.97 3382.33 3374.61 2742.49
-99% 4833.31 5749.98 6127.58 4967.29 4820.67 3711.22
-TABLE VIII: correctness rate/PRS area for each desired confidence level on fanshape distribution
-GMR trajectory predictor correctness rate
-confidence level NN BB(300) BB(6000) ILP(300) ILP(6000) optimal(30000)
-80% 93.1% 100% 100% 10.8% 3.9% 100%
-85% 95.2% 100% 100% 27.9% 16.2% 100%
-90% 99.8% 100% 100% 69.2% 55.8% 100%
-95% 95.6% 100% 100% 68.2% 65.4% 100%
-99% 100% 99.7% 100% 56.5% 75.8% 100%
-confidence level PRS area
-80% 2.75E-3 3.64E-3 3.46E-3 2.59E-3 2.39E-3 2.15E-3
-85% 3.44E-3 4.41E-3 4.35E-3 2.99E3 2.89E-3 2.39E-3
-90% 4.41E-3 7.37E-3 5.81E-3 3.83E-3 3.67E-3 2.99E-3
-95% 5.69E-3 7.89E-3 7.98E-3 5.03E-3 4.87E-3 3.95E-3
-99% 9.21E-3 11.19E-3 11.16E-3 8.01E-3 7.63E-3 6.46E-3
-TABLE IX: correctness rate/PRS area for each desired confidence level on GMR trajectory predictor distribution
-First, we aim to evaluate whether the proposed method can effectively produce accurate PRS.
-Table III presents the MSE and range comparisons between the proposed method and two baseline approaches, given the exact same visible points on the 2d binomial distribution.
-Our method consistently achieves lower MSE and a narrower range than the baseline methods, except at a 99% confidence level, where its performance closely matches that of ILP.
-Table III shows the MSE and range results on the fanshape distribution.
-The proposed method demonstrates exceptional performance, with significantly lower MSE and a tighter range across all confidence levels.
-Finally, Table III corroborates the findings from the previous tables, further showcasing that the proposed method achieves substantially lower MSE and a consistently tighter range, underscoring its superior performance.
-All in all, we found out that our method can produce very accurate PRS.
-All MSE is lower than 1%, with smaller intervals, indicating our method is very stable and consistent.
-In contrast, KDE-based baseline methods have higher MSE and variance especially when the desired confidence level is lower.
-On the other hand, it is not entirely fair to compare KDE-based methods directly, as they do not require a training process.
-Theoretically, estimating a distribution accurately with a small number of samples is nearly impossible.
-Therefore, it is also necessary to evaluate whether the trained neural networks can outperform the baseline methods when a sufficient number of samples are provided.
-Tabel VI VI VI present the MSE and range comparisons between the proposed method (can have only 100 visible samples) and two baseline methods, given 2000 visible samples on three test distributions.
-While the results of the two baseline methods show slight improvements, they still fall significantly short of the performance achieved by our method.
-Second, we want to ensure the convex bound generated by the neural networks is correct and useful.
-Table IX shows the performance of the neural network on 2d-binomial distribution.
-The neural network gets at least 98.7% correctness rate and only misses a few tests.
-Meanwhile, the average area we got is not significantly larger than the optimal area especially when $\alpha$ is 80%.
-As $\alpha$ increases, the difference between our area and the optimal area also increases.
-To guarantee reaching the confidence level, the neural networks tend to cover slightly more of the distribution.
-Since the distribution samples become sparser in the area far from the mean, even a small increase in coverage results in a significant increase in the area.
-Compared to the bounding box, while it may achieve a higher correctness rate due to its highly conservative nature, its area is significantly larger than that of the proposed method.
-Meanwhile, although the area of ILP is similar to that of NN, its correctness rate is significantly lower.
-Table IX shows the performance of the neural network on fanshape distribution.
-The neural network achieves a near-perfect correctness rate of close to 100%.
-Additionally, the area comparison follows a similar trend to that observed in the 2d binomial distribution.
-Table.
-IX shows the performance of the neural network on GMR trajectory predictor distribution.
-Although the correctness rate is not as high as in the previous tables, it still reaches a minimum of 93.1%.
-Meanwhile, the area remains superior to the Bounding Box, comparable to ILP, and only slightly larger than the optimal solution.
-In summary, the proposed neural network demonstrates consistently high correctness rates across various distributions, ranging from at least 93.1% to nearly 100%, while maintaining a compact area that is only marginally larger than the optimal solution.
-This balance between correctness and efficiency underscores the utility of the neural network-generated convex sets.
-The result supports that the Advantage 1 of the proposed method, NN can generate useful PRS with only a few samples.
-V-C Generating speed
-In addition to comparing the accuracy, we compare the generating time required to generate useful results.
+IV-B Motion Uncertainty
+For motion uncertainty, we evaluate flying obstacles subject to stochastic deviations in their trajectories.
+We simulate a flight using the Dubins vehicle model, where both speed and heading angle are sampled from a truncated Gaussian distribution.
+The resulting obstacle distribution for this simulated flight is shown in Fig.
+8a.
+[…]
+Correctness rates in Table III further highlight the robustness of our approach.
+NN achieves above 98% correctness at all CLs, while ILP often falls below 60% with limited samples.
+BB can achieve high correctness, but only at the cost of significantly larger probabilistic obstacle space areas, which reduces path planning efficiency.
+In contrast, NN produces obstacle spaces closer to the optimal size, especially at higher CLs.
+IV-C Behavior Uncertainty
+For behavior uncertainty, we evaluate flying obstacles with uncertain destinations.
+The real obstacle distribution is generated using a Gaussian Mixture Regression (GMR) trajectory predictor, which estimates the probability distribution of a flight’s future positions in terminal airspace [35].
+Given the past trajectory, the predictor models multiple possible future routes as a mixture of Gaussians.
+The resulting obstacle distribution for this predicted flight is shown in Fig.
+8b.
+Results under behavior uncertainty (Tables VIVI) reveal the strength of the learning-based approach.
+Despite the increased distributional complexity, NN maintains low MSE (below 2% for CL $\geq$ 85%) and narrow coverage ranges.
+By comparison, BB consistently overshoots the target coverage, while ILP suffers from high variance and instability, particularly with limited data.
+The correctness rates again demonstrate the advantage of NN.
+At 90% CL, NN achieves nearly perfect correctness (99.8%), whereas ILP with 300 samples drops to 69.2%.
+Importantly, NN generates obstacle spaces that are consistently smaller than BB while maintaining high coverage reliability.
+IV-D Sensor Uncertainty
+For sensor uncertainty, we evaluate obstacles whose positions are estimated from noisy measurements.
+In this scenario, the true positions of two adjacent obstacles are observed through a sensor corrupted by Gaussian noise, resulting in uncertainty about their exact locations.
+The real obstacle distribution is obtained by sampling noisy measurements of the true positions and constructing the corresponding probabilistic obstacle space.
+The resulting distribution is illustrated in Fig.
+8c.
+[…]
+Correctness rates remain above 98% for NN, ensuring that the generated obstacle spaces are reliable.
+Moreover, the obstacle space areas are significantly closer to the optimal solution compared with BB and ILP.
+This result keeps demonstrating NN’s ability to generalize well under noisy measurement conditions while avoiding excessive conservatism.
+IV-E Overall Discussion
+Across motion, behavior, and sensor uncertainty, the proposed NN consistently outperforms traditional BB and ILP methods. Three key observations can be drawn:
+[…]
+  • Reliability: NN maintains correctness rates above 93% in all scenarios, ensuring that the generated probabilistic obstacle spaces meet or exceed the desired confidence levels.
+  • Efficiency: NN produces compact obstacle spaces that are significantly smaller than BB while avoiding the instability of ILP.
+This efficiency is particularly critical in path planning applications where over-approximation reduces safe space.
+These results validate the effectiveness of the proposed learning-based framework in generating accurate and reliable probabilistic obstacle spaces under diverse forms of uncertainty.
+V Generating speed analysis
+Generating time is crucial in deciding whether a method can be used in online applications.
+Therefore, we compare the time to generate one correct probabilistic obstacle space by each method to prove our method is more useful in online applications.
 Fig.
@@
 The results clearly demonstrate that our method significantly outperforms the baseline approaches, achieving execution times that are an order of magnitude faster.
-The result supports that the Advantage 2 of the proposed method, NN can generate useful PRS very fast.
+The result supports that the Advantage 4 of the proposed method, NN can generate a useful probabilistic obstacle space very fast.
 Figure 9: Generating speed comparison
 VI Conclusion and Future
-In this paper, we propose using self-supervised learning to train a transformer-based neural network to generate convex approximations of probabilistic reachable sets.
-Once trained, the neural network can robustly generate convex approximations that cover the desired percentage of points sampled from the distribution of the probabilistic reachable set only given a few samples from the distribution.
+In this paper, we propose using supervised learning to train a transformer-based neural network to generate a confidence region of the real obstacle space distribution, which we define as the probabilistic obstacle space.
+Once trained, the neural network can robustly generate vertices of the polygon that cover the desired percentage of the real obstacle region, only given a few samples.
 The primary advantage of the proposed method is its extremely fast generation time while maintaining performance comparable to that of the baseline methods.
@@
 Only a few experts are required for the labels.
-No physical information about the system such as disturbance is required.
+No dynamic information about the obstacle such as disturbance is required.
 Acknowledgment
-The authors would like to acknowledge the support from the National Science Foundation under Grants CCF-2402689 and CMMI-2138612.
+The authors would like to acknowledge the support from the National Science Foundation under Grants CCF-2402689.
 Any opinions, findings, conclusions, or recommendations expressed in this paper are those of the authors and do not reflect the views of NSF.